### 11 Dec 2023 v2

# Diffusion Models for Reinforcement Learning: A Survey

Abstract
Diffusion models have emerged as a prominent class of generative models, surpassing previous methods regarding sample quality and training stability. Recent works have shown the advantages of diffusion models in improving reinforcement learning (RL) solutions, including as trajectory planners, expressive policy classes, data synthesizers, etc. This survey aims to provide an overview of the advancements in this emerging field and hopes to inspire new avenues of research. First, we examine several challenges encountered by current RL algorithms. Then, we present a taxonomy of existing methods based on the roles played by diffusion models in RL and explore how the existing challenges are addressed. We further outline successful applications of diffusion models in various RL-related tasks while discussing the limitations of current approaches. Finally, we conclude the survey and offer insights into future research directions, focusing on enhancing model performance and applying diffusion models to broader tasks. We are actively maintaining a GitHub repository for papers and other related resources in applying diffusion models in RL  [^1] https://github.com/apexrl/Diff4RLSurvey.
1 Introduction
Diffusion models have emerged as a powerful class of generative models, garnering significant attention in recent years. These models employ a denoising framework that can effectively reverse a multistep noising process to generate new data Song et al. (2021). In contrast to earlier generative models such as Variational Autoencoders (VAE) Kingma and Welling (2013) and Generative Adversarial Networks (GAN) Goodfellow et al. (2014), diffusion models exhibit superior capabilities in generating high-quality samples and demonstrate enhanced training stability. Consequently, they have made remarkable strides and achieved substantial success in diverse domains including computer vision Ho et al. (2020); Lugmayr et al. (2022); Luo and Hu (2021), natural language processing Austin et al. (2021); Li et al. (2022), audio generation Lee and Han (2021); Kong et al. (2020), and drug discovery Xu et al. (2022); Schneuing et al. (2022), etc.
Reinforcement learning (RL) Sutton and Barto (2018) focuses on training agents to solve sequential decision-making tasks by maximizing cumulative rewards. While RL has achieved remarkable successes in various domains Kober et al. (2013); Kiran et al. (2021), there are some long-standing challenges. Specifically, despite the considerable attention garnered by offline RL for overcoming low sample efficiency issue in online RL Kumar et al. (2020); Fujimoto and Gu (2021), conventional Gaussian policies may fail to fit the datasets with complex distributions for their restricted expressiveness. Meanwhile, although experience replay is used to improve sample efficiency Mnih et al. (2013), there is still data scarcity problem in environments with high-dimensional state spaces and complex interaction patterns. A common usage of learned dynamic models in model-based RL is planning in them Nagabandi et al. (2018); Schrittwieser et al. (2020); Zhu et al. (2021), but the per-step autoregressive planning approaches suffer from the compounding error problem Xiao et al. (2019). An ideal RL algorithm should be able to learn a single policy to perform multiple tasks and generalize to new environments Vithayathil Varghese and Mahmoud (2020); Beck et al. (2023). However, existing works still struggle in multitask generalizations.
Recently, there has been a series of works applying diffusion models in sequential decision-making tasks, with a particular focus on offline RL. As a representative work, Diffuser Janner et al. (2022) fits a diffusion model for trajectory generation on the offline dataset, and plans desired future trajectories by guided sampling. There have been many following works where diffusion models behave as different modules in the RL pipeline, e.g., replacing conventional Gaussian policies Wang et al. (2023), augmenting experience dataset Lu et al. (2023b), extracting latent skills Venkatraman et al. (2023), among others. We also observe that planning and decision-making algorithms facilitated by diffusion models perform well in broader applications such as multitask RL He et al. (2023a), imitation learning Hegde et al. (2023), and trajectory generation Zhang et al. (2022). More importantly, diffusion models have already shed light on resolving those long-standing challenges in RL owing to their powerful and flexible distributional modeling ability.
This survey centers its attention on the utilization of diffusion models in RL, with additional consideration given to methods incorporating diffusion models in the contexts of trajectory generation and imitation learning, primarily due to the evident interrelations between these fields. Section 2 elaborates on the aforementioned RL challenges, and discusses how diffusion models can help solve each challenge. Section 3 provides a background on the foundations of diffusion models and also covers two class of methods that are particularly important in RL-related applications: guided sampling and fast sampling. Section 4 illustrates what roles diffusion models play in RL among existing works. Section 5 discusses the contribution of diffusion models on different RL-related applications. In Section 6, we point out the limitations when applying diffusion models and compare them with the transformer-based methods. Section 7 summarizes the survey with a discussion on emerging new topics.
2 Challenges in Reinforcement Learning
Figure 1: An illustration of how diffusion models play a different role in the classic Agent-Environment-Buffer cycle compared to previous solutions. (1) When used as a planner, diffusion models optimize the whole trajectory at each denoising step, whereas the autoregressive models generate the next-step output only based on previously planned partial subsequences. (2) When used as a policy, diffusion models can model arbitrary action distributions, whereas Gaussian policies can only fit the possibly diversified dataset distribution with unimodal distributions. (3) When used as a data synthesizer, diffusion models augment the dataset with generated data sampled from the learned dataset distribution, whereas augmentation with random perturbations might generate samples that deviate from data samples.
In this section, we list four challenges in RL algorithms and briefly discuss why diffusion models can address them.
2.1 Restricted Expressiveness in Offline Learning
Online RL Sutton and Barto (2018); Arulkumaran et al. (2017) has been criticized for low sample efficiency, which makes it difficult to be applied in real-world scenarios. Offline RL Fujimoto et al. (2019); Kumar et al. (2020); Fujimoto and Gu (2021), which learns optimal policies purely from pre-collected datasets, obviates the need to interact with the environment during training, and can significantly improve sample efficiency. Directly applying off-policy RL methods Lillicrap et al. (2015); Haarnoja et al. (2018) to offline learning suffers from the extrapolation error problem Fujimoto et al. (2019). Existing works either penalize the value predictions on out-of-distribution samples Kumar et al. (2020); Nachum et al. (2019), or constrain the learning policy to be close to the data collection policy Wu et al. (2019); Kostrikov et al. (2021). However, penalties on the value function can make the learned policy over-conservative Lyu et al. (2022); for algorithms using policy constraints, since the policy is usually parameterized as a unimodal Gaussian, the restricted expressiveness makes it hard to fit the possibly diversified dataset. The reinforcement learning via supervised learning framework (RvS) Schmidhuber (2019); Srivastava et al. (2019) is now becoming another essential paradigm in offline RL, which bypasses the need for Q learning and is thus free of extrapolation errors. RvS learns a policy conditioned on the observed returns via supervised learning and then conditions the learned policy on a high enough return during online evaluations to generate desired behaviors Chen et al. (2021); Lee et al. (2022). Similar to policy constraining methods, RvS requires fitting the entire offline dataset. Therefore, the expressiveness of parametrized policies also matters in RvS. Since diffusion models can represent arbitrary normalizable distributions Neal and others (2011), they hold potential to effectively improve the performance of policy constraining and RvS algorithms on complex datasets.
2.2 Data Scarcity in Experience Replay
Off-policy and offline RL methods use different levels of experience replay to improve sample efficiency. Note that experience replay in some cases only refers to data reuse in off-policy RL. Here, we use the term to broadly refer to updating the current model with rollout data from other policies. In off-policy RL, although all previously collected experiences can be used for policy learning, the available data during training may still be inadequate for effective policy optimization due to the speed limit of simulations and the potentially large state and action spaces. In offline RL, as no further interactions are allowed, policy learning is more limited by the quality and coverage of the dataset. Inspired by the success of data augmentation in computer vision, some works implement similar augmentation techniques in RL to mitigate the data scarcity problem. RAD Laskin et al. (2020) uses typical image augmentation techniques such as random cropping or rotation to boost the learning efficiency in vision-based RL. Imre (2021) and Cho et al. (2022) use generative models, VAE Kingma and Welling (2013) and GAN Goodfellow et al. (2014), to augment the real dataset with synthetic data sampled from the learned data distribution. However, existing works either suffer from a lack of fidelity when using random augmentation or are limited to simple environments due to insufficient modeling ability of particular generative models, making it difficult for them to extend to more complex tasks. Diffusion models have already demonstrated notable performances surpassing previous generative models in domains including image and video synthesis Ho et al. (2020, 2022). When applied to RL data, diffusion models are more suitable for augmenting high-dimensional datasets with intricate interactions.
2.3 Compounding Error in Model-based Planning
Model-based RL (MBRL) Luo et al. (2022); Moerland et al. (2023) fits a model of dynamic transitions either from data obtained from online rollout or an offline dataset, and expects the model to facilitate decision-making. Common dynamic models mimic single-step state transitions and rewards in the dataset. Due to the limited data support and the possible stochasticity of the ground-truth transition function, there could be single-step errors when predicting with a neural network model. As a result, the compounding error problem arises when using the model for multistep planning Xiao et al. (2019), i.e., cumulative single-step errors can cause the planned states to deviate from the dataset distribution, further increasing the error in the subsequent model predictions. In contrast, diffusion models with powerful modeling ability of joint distributions can operate on the trajectory level and plan for all time steps simultaneously, offering better temporal consistency and less compounding errors.
2.4 Generalization in Multitask Learning
Normal RL algorithms lack generalization abilities at the task level Vithayathil Varghese and Mahmoud (2020). Even in the same environment, changing the reward function requires retraining a policy from scratch. Existing works studying online multitask RL Yu et al. (2020); Liu et al. (2021) attempt to learn the same policy on different task environments, suffering from the problem of conflicting gradients across multiple tasks, as well as low sample efficiency due to pure online learning. Recently, it has become a trending research direction to utilize a high-capacity model trained on multitask offline datasets and then deployed on new tasks with or without online fine-tuning Taiga et al. (2022); Oh et al. (2017). Transformer-based pre-training decision models like Gato Reed et al. (2022) achieve notable successes in multitask policy learning. However, they typically require high-quality datasets and come with large parameter sizes as well as corresponding high training and inference costs. How to design an algorithm that can efficiently fit multitask datasets of mixed quality and generalize to new tasks emerges as a vital issue in multitask RL. As a powerful class of generative models, diffusion models can handle multimodal distributions in multitask datasets, and adapt to new tasks by estimating the task distribution.
3 Foundations of Diffusion Models
This section provides the foundations of diffusion models. Two prominent formulations are presented: Denoising Diffusion Probabilistic Model (DDPM) Ho et al. (2020) and Score-based Generative Models Song et al. (2021). DDPM is widely used due to its simplicity, whereas the score-based formulation extends it to encompass continuous-time diffusion processes. Furthermore, guided sampling methods play a critical role in integrating diffusion models into RL frameworks. These methods can be divided into two main categories based on their approach to guiding the sampling process: classifier-guidance Dhariwal and Nichol (2021), which requires an extra classifier, and classifier-free guidance Ho and Salimans (2022), which makes guiding conditions as part of the model input. Additionally, in order to boost sampling speed especially during online interactions, fast sampling techniques have been adopted when using diffusion models in RL-related tasks Kang et al. (2023); Chi et al. (2023). Here we briefly cover some representative works in studying fasting sampling of diffusion models, including learning-based and learning-free methods.
3.1 Denoising Diffusion Probabilistic Model
Assuming that the real data, denoted as $x^{0}$, is sampled from an underlying distribution $q(x^{0})$, the Denoising Diffusion Probabilistic Model (DDPM) utilizes a parameterized diffusion process, represented as $p_{\theta}(x^{0})=\int p(x^{T})\prod_{t=1}^{T}p_{\theta}(x^{t-1}|x^{t})\mathop{}\!\mathrm{d}x^{1:T}$, to model how the pure noise, denoted as $x^{T}=\mathcal{N}(0,I)$, is denoised into real data $x^{0}$. In this formulation, each step of the diffusion process is represented by $x^{t}$, with $T$ indicating the total number of steps. It is important to note that both diffusion models and RL use time step notations; thus, following common practice in RL, we denote diffusion steps with superscripts and RL time steps with subscripts. The sequence $x^{T:0}$ is defined as a Markov chain with learned Gaussian transitions characterized by
$p_{\theta}(x^{t-1}|x^{t})=\mathcal{N}(\mu_{\theta}(x^{t},t),\Sigma(x^{t},t))~{}.$
If the process is reversed to $x^{0:T}$, each step is the forward transition $q(x^{t}|x^{t-1})$. This transition can be interpreted as adding Gaussian noise to the data according to a variance schedule $\beta^{1:T}$:
$x^{t}=\sqrt{\alpha^{t}}x^{t-1}+\sqrt{1-\alpha^{t}}\epsilon_{t}~{},$ (1)
where $\alpha^{t}=1-\beta^{t}$, $\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. $q$ also holds the Markov property as $q(x^{t}|x^{t-1})$ only depends on $x^{t-1}$. In general, $\alpha^{T}$ should satisfy $\lim_{T\rightarrow+\infty}\alpha^{T}\rightarrow 0$ to ensure that $x^{T}$ would be a pure standard Gaussian noise when $T$ is sufficiently large. And from Eq. (1), we can infer that
$x^{t}=\sqrt{\bar{\alpha}^{t}}x^{0}+\sqrt{1-\bar{\alpha}^{t}}{\epsilon(x^{t},t)}~{},$ (2)
where $\bar{\alpha}^{t}=\prod_{1}^{t}\alpha^{i}$ and ${\epsilon(x^{t},t)}\sim\mathcal{N}(0,\mathbf{I})$ is the integrated noise from step 0 to $t$. $\epsilon(x,t)$ is unknown and to be learned by a network. So far, we can calculate the distribution from step 0 to step $t$ by simply one calculation. As for the denoising process $q(x^{t-1}|x^{t})$, it is complex while $q(x^{t-1}|x^{t},x^{0})$ is relatively simpler. Based on Bayes Theorem and the Markov property of $q$, we have
$q(x^{t-1}|x^{0},x^{t})=\frac{q(x^{t-1}|x^{0})}{q(x^{t}|x^{0})}q(x^{t}|x^{t-1})=\mathcal{N}(\mu,\sigma^{2})~{},$
where
$\sigma^{2}(x^{t},t)$ $=\beta^{t}\frac{1-\bar{\alpha}^{t-1}}{1-\bar{\alpha}^{t}},$ (3)
$\mu(x^{t},t)$ $=\frac{1-\bar{\alpha}^{t-1}}{1-\bar{\alpha}^{t}}\sqrt{\alpha^{t}}x^{t}+\frac{\sqrt{\bar{\alpha}^{t-1}}\beta^{t}}{1-\bar{\alpha}^{t}}x^{0}~{}.$
From Eq. (2), we can eliminate $x^{0}$, yielding:
$\mu(x^{t},t)=\frac{1}{\sqrt{{\alpha}^{t}}}(x^{t}-\frac{\beta^{t}}{\sqrt{1-\bar{\alpha}^{t}}}\epsilon(x^{t},t))~{}.$ (4)
This allows us to sample $x^{T}$ from standard Gaussian noise and progressively denoise it step by step until we obtain $x^{0}$. However, the noise variable $\epsilon$ is still unknown. To address this, a network $\epsilon_{\theta}$, parameterized by $\theta$, is employed to learn the noise generation process. In general, the loss function is the distance between the generated $\mu(x^{t},t)$ using random variables $\epsilon$ and the network output $\epsilon_{\theta}$. Specially, $\mu(x^{t},t)$ is defined as in Eq. (4), and $\mu_{\theta}(x^{t},t)=\frac{1}{\sqrt{{\alpha}^{t}}}(x^{t}-\frac{\beta^{t}}{\sqrt{1-\bar{\alpha}^{t}}}\epsilon_{\theta}(x^{t},t))$.
Ho et al. (2020) discover that the following version of the loss function ignoring the weights has a better performance in experiments:
$\mathcal{L}_{t-1}\coloneqq\mathbb{E}_{x^{0},\epsilon}[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}^{t}}x^{0}+(\sqrt{1-\bar{\alpha}^{t}})\epsilon,t)\|^{2}]~{}.$ (5)
Here, $\epsilon$ is sampled from the standard Gaussian distribution $\mathcal{N}(0,\mathbf{I})$ and $\mathcal{L}_{t-1}$ represents the loss function at diffusion step $t-1$, when $t>1$.
3.2 Score-based Generative Models
DDPM illustrates its iterative sampling process in a step-by-step manner, with each step taking place at a discrete time interval. Song et al. (2021) extend DDPM to continuous-time cases, where the sequence $x^{0},x^{1},\ldots,x^{T}$ is replaced with a continual function $x^{t},t\in[0,T]$. Now, the forwarding process can be described as a Stochastic Differential Equation (SDE):
$\mathop{}\!\mathrm{d}x=f(x,t)\mathop{}\!\mathrm{d}t+g(t)\mathop{}\!\mathrm{d}w~{},$
where $f(x,t)$ and $g(t)$ are pre-defined functions, and $\mathop{}\!\mathrm{d}w$ is the Brownian Motion. The example DDPM process in Eq. (1) is represented as $f(x,t)=-\frac{1}{2}\beta(t)x$, $g(t)=\sqrt{\beta(t)}~{}.$ According to Langevin Dynamics, the reverse of the forwarding process, as the sampling process, is described by a reverse-time SDE:
$\mathop{}\!\mathrm{d}x=[f(x,t)-g^{2}(t)\nabla_{x}\log q_{t}(x)]\mathop{}\!\mathrm{d}t+g(t)\mathop{}\!\mathrm{d}\bar{w}~{},$ (6)
where $\bar{w}$ is the reverse Brown Motion. The gradient term $\nabla_{x}\log q_{t}(x)$ is called the score function of the distribution $q_{t}(x)$. As the score function remains unknown, score matching Song et al. (2019) proposes that a parameterized score-based model $\epsilon_{\theta}(x,t)$ can be trained to approximate it. In practice, diffusion models usually match the scaled score function $-\sigma_{t}\nabla_{x}\log q_{t}(x)$. Thus, a time-dependent score model can be trained by minimizing the Fisher Divergence:
$\min_{\theta}\mathbb{E}_{x^{0}\sim q_{0},x^{t}\sim q_{0t}}\left[\|\epsilon_{\theta}(x^{t},t)+\sigma_{t}\nabla_{x^{t}}\log q_{0t}(x^{t}|x^{0})\|_{2}^{2}\right]~{}.$
Solving the SDE in Eq. (6) is difficult even with a trained score model. Alternatively, we can solve the corresponding Probability Flow ODE (PF-ODE), whose marginal distribution at any time is equal to that given by the original SDE in Eq. (6):
$\mathop{}\!\mathrm{d}x=\left[f(x,t)-\frac{1}{2}g^{2}(t)\nabla_{x}\log q_{t}(x)\right]\mathop{}\!\mathrm{d}t~{}.$ (7)
Given a trained score model, all terms in Eq. (7) are known thus it can be efficiently solved by various ODE solvers such as VODE Brown et al. (1989). In this case, a black-box ODE solver Dormand and Prince (1980) is highly recommended because it not only produces high-quality samples but also allows us to explicitly trade-off accuracy for efficiency.
3.3 Guided Sampling Methods
Diffusion models with guided sampling methods care about the conditioned data distribution $p(x|y)$, which makes it possible to generate samples with attributes of the label $y$. Based on whether an extra classifier model is to be trained, the methods are divided into two categories: classifier guidance and classifier-free guidance. An advantage of classifier guidance sampling is that the classifier and the diffusion model are trained independently. If you already have a diffusion model, you just need to train a classifier and integrate it with your diffusion model while sampling. Classifier-free guidance retrains the model totally, but it is more expressive and performs better compared to classifier guidance sampling methods.
3.3.1 Classifier Guidance
A simple thought is to train a differentiable discriminative model that presents $p(y|x)$. In other words, all we need to do is to train an extra classifier $p(y|x^{t})$ trained on noisy samples $x^{t}$. Assume the classifier has been pre-trained. Then, the reversing process is described as
$p_{\theta,\phi}(x^{t}|x^{t+1},y)=Zp_{\theta}(x^{t}|x^{t+1})p_{\phi}(y|x^{t})~{},$ (8)
where $Z$ is a normalization factor. Dhariwal and Nichol (2021) state that Eq. (8) can be approximately regarded as another Gaussian distribution:
$p(x^{t}|x^{t+1},y)=\mathcal{N}(\mu(x^{t},t)+s\Sigma(x^{t},t)g,\Sigma(x^{t},t))~{},$
where $g=\nabla_{x^{t}}\log p_{\phi}(y|x^{t})|_{x^{t}=\mu}$ and $s$ is the guidance scale to control the effect of conditions. The $\mu(x^{t},t)$ and $\Sigma(x^{t},t)$ function is identical to those in DDPM (Eq. (4) and Eq. (3)). Therefore, the sampling process can be explicitly presented as follows:
$x^{t+1}\sim\mathcal{N}(\mu+s\Sigma g,\Sigma)~{},$
where $\Sigma$ is short for $\Sigma(t,x^{t})$.
3.3.2 Classifier-free Guidance
Rather than estimating the conditioned data distribution $p(x|y)$ directly, classifier-free methods try to predict the score function $\nabla_{x}\log p(x|y)$. With the Bayes Theorem, we can decompose the score function into an unconditional term and a classifier conditioning term as
$\nabla_{x}\log p(x|y)=\nabla_{x}\log p(y|x)+\nabla_{x}\log p(x)~{}.$ (9)
Unlike the classifier guidance, the original training setup is modified so the diffusion model should be retrained. The network to estimate noise is defined as $\epsilon(x^{t},c)$ ($c=\varnothing$ for unconditional inputs). In practice, the conditional and unconditional models are trained with the same set of network parameters by randomly setting $c=\varnothing$ with a pre-specified probability during training. From Eq. (9) we can infer
$\nabla_{x}\log p(y|x)=\nabla_{x}\log p(x|y)-\nabla_{x}\log p(x)~{}.$
As Song et al. (2022) reveal, diffusion models and the score functions are equivalent, which indicates $\nabla_{x^{t}}\log p(x^{t})\propto\epsilon(x^{t},t)$. As a result, we can now substitute $\nabla_{x}\log p(y|x)$ with this into the formula for classifier guidance
$\bar{\epsilon}_{w}(x^{t},y)$ $=\epsilon_{\theta}(x^{t},y)+w(\epsilon_{\theta}(x^{t},y)-\epsilon_{\theta}(x^{t}))$
$=(1+w)\epsilon_{\theta}(x^{t},y)-w\epsilon_{\theta}(x^{t})~{},$
where $w$ stands for the guidance scale.
3.4 Fast Sampling Methods
Diffusion model has been criticized for its prolonged iterative sampling time. Several fast sampling skills are proposed to solve this issue. Generally, the skills extend the diffusion model to a more fundamental paradigm to acquire efficiency. We categorize the methods into two categories: those that do not involve learning (learning-free) and those that require extra learning sessions (learning-based).
Learning-free sampling methods.
Denoising Diffusion Implicit Models (DDIM) Song et al. (2022) is one of the earliest works on sampling acceleration. It aims at extending DDPM to a non-Markovian case by learning another Markov chain $q_{\theta}(x^{t-1}|x^{t},x^{0})$. The work also reveals that DDPM and DDIM are special cases of a more general paradigm. Moreover, it is observed in later works that DDIM is the discrete version of solving the PF-ODE Eq. (7). Then, some high-order solvers emerged, such as DPM-solver Lu et al. (2022), which provides an excellent trade-off between sample quality and sampling speed. With DDIM as its first-order version, DPM-solver boosts the efficiency of solving PF-ODE, outperforming common numerical ODE solvers like Runge-Kutta. Consequently, DPM-solver has become one of the most frequently used fast sampling methods.
Learning-based sampling methods.
Learning-based sampling is another approach to fast sampling. Unlike the learning-free methods, they include extra learning processes to reach a higher sampling efficiency at a slight expense of sampling quality. A recent work, Truncate Diffusion Probabilistic Model (TDPM) Zheng et al. (2023), proposes that both the diffusion and denoising process can be truncated so that the iterative steps are reduced. Specifically, the forward process is truncated when the sample is noisy enough, and the denoising process starts from a relatively noisy sample (not pure noise), which will be learned by a network or else. Moreover, Watson et al. (2021) learn a strategy to select the best $K$ time steps to maximize the training objective for the DDPMs, which also decreases the denoising steps.
Table 1: Summary of papers on diffusion models for RL.
Model & Paper Role of Diffusion Models Keyword(s) Guidance
Diffuser Janner et al. (2022) Planner Offline Classifier
AdaptDiffuser Liang et al. (2023) Offline Classifier
EDGI Brehmer et al. (2023) Offline Classifier
TCD Hu et al. (2023) Offline Classifier-free
Crossway Diffusion Li et al. (2023b) Offline; Robotics None
SGP Suh et al. (2023) Offline; Robotics None
GSC Mishra et al. (2023) Imitation; Robotics Classifier
UniPi Du et al. (2023a) Imitation; Robotics Classifier-free
ChainedDiffuser Xian et al. (2023) Imitation; Robotics None
Diffusion Policy Chi et al. (2023) Imitation; Robotics None
AVDC Ko et al. (2023) Imitation; Robotics None
MTDiff-p He et al. (2023a) Offline; Multi-task Classifier-free
MetaDiffuser Ni et al. (2023) Offline; Multi-task Classifier-free
SafeDiffuser Xiao et al. (2023) Offline; Safe Self-proposed
MADiff Zhu et al. (2023) Offline; Multi-agent Classifier-free
HDMI Li et al. (2023a) Offline; Hierarchical Classifier-free
MLD Chen et al. (2022) Trajectory Generation Classifier-free
MDM Tevet et al. (2022) Trajectory Generation Classifier-free
UniSim Yang et al. (2023) Trajectory Generation Classifier-free
ReMoDiffuse Zhang et al. (2023b) Trajectory Generation Classifier-free
SinMDM Raab et al. (2023) Trajectory Generation None
EquiDiff Chen et al. (2023b) Trajectory Generation None
MoFusion Dabral et al. (2022) Trajectory Generation None
MotionDiffuse Zhang et al. (2022) Trajectory Generation None
MPD Carvalho et al. (2023) Trajectory Generation; Robotics Classifier
MotionDiffuser Jiang et al. (2023) Trajectory Generation; Multi-agent Classifier
AlignDiff Dong et al. (2023) RLHF Classifier-free
Diffusion-QL Wang et al. (2023) Policy Offline Q-loss
SRDP Ada et al. (2023) Offline Q-loss
SfBC Chen et al. (2023a) Offline Policy Gradient
IDQL Hansen-Estruch et al. (2023) Offline Policy Gradient
EDP Kang et al. (2023) Offline Policy Gradient
DiffCPS He et al. (2023b) Offline None
NoMaD Sridhar et al. (2023) Robotics None
BESO Reuss et al. (2023) Offline; Goal-conditioned Classifier-free
CEP Lu et al. (2023a) Offline; Image Synthesis Policy Gradient
DOM2 Li et al. (2023c) Offline; Multi-agent Q-loss
CPQL Chen et al. (2023d) Offline; Online Q-loss
 Pearce et al. (2023) Imitation Classifier-free
 Yoneda et al. (2023) Imitation; Robotics None
PlayFusion Chen et al. (2023c) Imitation; Robotics None
XSkill Xu et al. (2023) Imitation; Robotics None
CoDP Ng et al. (2023) Human-in-the-loop None
GenAug Chen et al. (2023e) Data Synthesizer Robotics None
ROSIE Yu et al. (2023) Robotics None
SynthER Lu et al. (2023b) Offline; Online None
MTDiff-s He et al. (2023a) Offline; Multi-task Classifier-free
LDCQ Venkatraman et al. (2023) Latent Representation Offline Classifier-free
LatentDiffuser Li (2023) Offline Policy Gradient
 Hegde et al. (2023) Quality Diversity None
DVF Mazoure et al. (2023) Value Function Offline None
4 The Roles of Diffusion Models in RL
Figure 2: Different roles of diffusion models in RL. (a) Diffusion models as the planner. The sampling target is a part of trajectories whose components may vary from specific tasks. (b) Diffusion models as the policy. The sampling target is the action conditioned on the state, usually guided by the Q-function via policy gradient-style guidance or directly subtracting it from the training objective. (c) Diffusion models as the data synthesizer. The sampling target is also the trajectory, and both real and synthetic data are used for downstream policy improvement. For better visualizations, we omit the arrows for $N$ denoising iterations in (c) and only show generated synthetic data from randomly sampled noise. Note that there are other roles that are less explored, and we introduce them in Section 4.4.
Diffusion models have proven their ability to generate diverse data and model multi-modal distributions. Considering the long-existing challenges introduced in Section 2, it is sufficient to improve the performance and sample efficiency of RL algorithms with diffusion models. In Fig. 1, we illustrate how diffusion models play a different role in RL compared to previous solutions. Current works applying diffusion models on RL mainly fall into three categories: using diffusion models as the planner, as the policy, and as the data synthesizer. It is essential to note that we include methods that generate action-only sequences as planners, even though some of the representative works have “policy” in their names, e.g., Diffusion Policy Chi et al. (2023). Generating multi-step action sequences can be viewed as planning in action space, and the use of diffusion models to ensure temporal consistency is similar to other planning-based diffusion methods. The following subsections will illustrate overall frameworks and representative papers for each category.
4.1 Planner
Planning in RL refers to the process of using a model of the environment to make decisions imaginarily, and then selecting the best action in order to maximize a cumulative reward signal. This process usually simulates or explores different sequences of actions and states, predicting the outcomes of its decisions, thus resulting in better actions from the perspective of a longer horizon. Therefore, planning is commonly applied in the MBRL framework. However, the decision sequences used for planning are generated autoregressively, which may lead to severe compounding errors, especially in the offline setting, due to the limited data support. Diffusion models provide a possible solution since they can generate the whole sequence simultaneously.
A general framework of diffusion models as planners is shown in Fig. 2(a), which was first proposed by Janner et al. (2022). Inputs and outputs of the diffusion model are usually clips of the real trajectory $\tau=(s_{1},a_{1},r_{1},\ldots,s_{H},a_{H},r_{H})$, denoted as $x(\tau)=(e_{1},e_{2},\ldots,e_{H})$. Here $e_{t}$ represents for the selected elements from $(s_{t},a_{t},r_{t})$, where various choices can be made as $e_{t}=(s_{t},a_{t})$ Janner et al. (2022); Liang et al. (2023); He et al. (2023a); Xiao et al. (2023), or $e_{t}=(s_{t},a_{t},r_{t})$ He et al. (2023a); Hu et al. (2023), $e_{t}=s_{t}$ Ajay et al. (2023); Zhu et al. (2023), or $e_{t}=a_{t}$ Chi et al. (2023); Li et al. (2023b). $y(\tau)$ is the guidance that contains the desired property of the generated trajectory, such as the discounted return, the indicator of whether the task is completed, etc. Such guidance is meant to lead the diffusion model to generate good trajectories instead of those that only satisfy the environment dynamics. Suppose the diffusion model $\epsilon_{\theta}$ parameterized by $\theta$ predicts the noise added in the forward process, and $x^{k}\coloneqq x(\tau^{k})$ is the corresponding clip of $\tau$ at the $k$-th diffusion step. Either classifier guidance or classifier-free guidance is incorporated. For classifier guidance Janner et al. (2022); Liang et al. (2023), $\epsilon_{\theta}$ is trained as usual by taking $\tau^{k}$ and the timestep $k$ as input:
$\mathcal{L}(\theta)=\mathbb{E}_{\tau,k,\epsilon}[\|\epsilon-\epsilon_{\theta}(x^{k},k)\|_{2}^{2}]~{}.$
After training, suppose $\mu_{\theta}(x^{k},k)$ and $\Sigma$ are the mean and the variance of the original reverse process $p_{\theta}(x^{k-1}|x^{k})$, respectively. The gradient of $y(\tau)$ is injected during sampling as the guidance:
$p_{\theta}(x^{k-1}|x^{k},y(\tau))\approx\mathcal{N}(x^{k-1};\mu_{\theta}+\alpha\Sigma^{k}\nabla y(\tau),\Sigma^{k})~{},$
where $\alpha$ is a hyperparameter. In contrast, classifier-free guidance Ajay et al. (2023); He et al. (2023a); Hu et al. (2023) injects $y(\tau)$ in both training and sampling stages. The noise prediction model $\epsilon_{\theta}$ now takes an additional $y(\tau)$ as input, giving the training objective:
$\mathcal{L}(\theta)=\mathbb{E}_{\tau,k,\epsilon,\beta}[\|\epsilon-\epsilon_{\theta}(x^{k},(1-\beta)y(\tau)+\beta\varnothing,k)\|_{2}^{2}]~{},$
where $\beta$ is sampled from a Bernoulli distribution with probability $p$. At sampling, $y(\tau)$ is directly fed into $\epsilon_{\theta}$, as
$p_{\theta}(x^{k-1}|x^{k},y(\tau))\approx\mathcal{N}(x^{k-1};\mu_{\theta}(x^{k},y(\tau),k),\Sigma^{k})~{}.$
As the time consumption of sampling using the original DDPM is unacceptable when deploying RL in the online environment, fast sampling methods are commonly incorporated, and the diffusion step for sampling is usually set to 100 or 200. Besides, to guarantee sample stability during the sampling process, the first several steps of a sampled trajectory can be replaced with the ground truth as a hard constraint. For example, when planning the trajectory after $e_{1:h}$, the model first generates the full trajectory $\tilde{x}^{k}=\tilde{e}_{1:H}^{k}$ at the $k$-the diffusion step. Then the first $h$ steps are replaced, forming $\tilde{x}^{k}\leftarrow(e_{1},e_{2},\ldots,e_{h},\tilde{e}_{h+1}^{k},\ldots,\tilde{e}_{H}^{k})$.
The core ideas behind why non-autoregressive diffusion models can generate Markovian decision sequences are stated by Janner et al. (2022). With the U-Net backbone, the sampled token $e_{t}^{k}$ takes information in a local receptive field $e_{t-l:t+l}^{k+1}$, where $l$ is determined by the model structure. Once the local consistency is guaranteed, the global consistency will also be built during the denoising process as information spreads among the trajectory. Though this raises new confusion that using future information is anti-causal for sequential problems, RL includes a hidden hypothesis that the current decision will lead to optimal future outcomes. That is, utilizing future information does not violate the general RL framework.
Benefiting from the non-autoregressive scheme, diffusion models surpass traditional models in generating planning sequences. First, such generations do not suffer from the aforementioned compounding errors. Then, diffusion models fit for both continuous-reward and sparse-reward settings Janner et al. (2022), even for more challenging multitask planning He et al. (2023a), safe planning Xiao et al. (2023), and generating based on multi-modal observations Du et al. (2023a); Chi et al. (2023). Besides, diffusion models can also learn from non-Markovian data by exploiting temporal information Hu et al. (2023).
4.2 Policy
Compared with traditional RL taxonomy, which roughly divides RL algorithms into MBRL and model-free RL, using diffusion models as the planner is similar to MBRL and focuses on capturing the environment dynamics. In contrast, taking diffusion models as the policy follows the framework of model-free RL. Section 2.1 states the main drawbacks of offline policy learning frameworks: over-conservatism and poor capability on diversified datasets. With excellent expressive ability on multi-modal distribution, many works utilize diffusion models as the policy to tackle these problems.
Diffusion-QL Wang et al. (2023) first explores the advantages of the diffusion policy in offline RL and finds that it can perfectly fit on dataset generated by strong multi-modal behavior policies, where previous distance-based BC methods fail. Experiment results also support the dominance of diffusion policies on diversified datasets Lu et al. (2023a); Chen et al. (2023a); Hansen-Estruch et al. (2023); Ada et al. (2023); Li et al. (2023c). Compared with diffusion models as planners, the diffusion target of the diffusion policy is simply the action given the current state, as shown in Fig. 2(b). Suppose the noise predictor is $\epsilon_{\theta}(a^{k},k,s)$ parameterized by $\theta$, and the derived diffusion model is $\mu_{\theta}(a|s)$. To guide the model sampling actions with high Q-values, it is necessary to take $Q(s,a)$ into consideration. Some papers Chen et al. (2023a); Lu et al. (2023a); Hansen-Estruch et al. (2023); Kang et al. (2023) construct the policy by (advantage) weighted regression as
$\pi_{\theta}(a|s)\propto\mu_{\theta}(a|s)\exp(\alpha Q(s,a))~{},$
where $\alpha$ is a hyperparameter. Following this, Chen et al. (2023a) decouple the policy learning into behavior learning and action evaluation, which allows more freedom in the choice of guidance. It also proposes in-sample planning for Q-learning, which avoids extrapolation errors in previous offline RL methods. Lu et al. (2023a) further generalize this framework to any energy reweighted distribution $p(x)\propto q(x)\exp(-\beta\mathcal{E}(x))$, where $\mathcal{E}(x)$ is an energy function, and train it via contrastive learning. Approaches using explicit Q-values as guidance are also considered. Wang et al. (2023) and Ada et al. (2023) subtract a weighted expectation of $Q(s,a)$ in the training loss as
$\mathcal{L}(\theta)$ $=\mathbb{E}_{k,\epsilon,(s,a)\sim\mathcal{D}}[\|\epsilon-\epsilon_{\theta}(a^{k},s,k)\|_{2}^{2}]$
$\quad-\frac{\eta}{\mathbb{E}_{(s,a)\sim\mathcal{D}}[Q(s,a)]}\cdot\mathbb{E}_{s\sim\mathcal{D},a^{0}\sim\pi_{\theta}(\cdot|s)}[Q(s,a^{0})]~{},$
where $\eta$ is a hyperparameter, and $\mathcal{D}$ is the offline dataset.
Fast reaction is crucial when deploying policies in online environments. Therefore, almost all diffusion policies use smaller diffusion steps during sampling, usually about 15 steps. ODE solvers such as the DPM-solver Lu et al. (2022) are also used to accelerate sampling Chen et al. (2023a); Lu et al. (2023a); Kang et al. (2023); Li et al. (2023c)Pearce et al. (2023) compare three different backbone structures of diffusion models, finding that MLP-sieve or Transformer is also sufficient in expressiveness. Kang et al. (2023) introduce action approximation, which allows one-step action sampling in the training stage.
4.3 Data Synthesizer
In addition to fitting multi-modal distributions, a simple and common use of diffusion models is to generate more training samples, which has been widely applied and proven in computer vision. Therefore, it is natural to apply the diffusion model as a data synthesizer on RL datasets, as data scarcity is a practical challenge of RL, as stated in Section 2.2. To guarantee consistency of synthetic data to the environment dynamics, previous data augmentation approaches in RL usually add small perturbations to existing states and actions Sinha et al. (2021). In contrast, Fig. 2(c) illustrates that diffusion models learn the data distribution from the whole dataset $\mathcal{D}$, and enable generating highly diversified data while keeping consistency. Lu et al. (2023b) investigate the ability of diffusion models as the data synthesizer in both offline and online settings. It directly trains the diffusion model from the offline dataset or the online replay buffer and then generates more samples for policy improvement. Analysis shows that the quality of data generated by diffusion models is higher than those generated by explicit data augmentation in diversity and accuracy. With synthetic data, the performance of the offline policy and the sample efficiency of the online policy are significantly improved. He et al. (2023a) deploy diffusion models to augment data for multitask offline datasets and achieves better performance than those on single-task datasets. It claims that fitting on multiple tasks may enable implicit knowledge sharing across tasks, which also benefits from the multi-modal property of diffusion models.
4.4 Others
Besides directions that have been mainly focused on, some ways of improving RL with diffusion models are under exploration. Hegde et al. (2023) take a similar idea as hyper networks in meta-learning, generating parameters of policies for quality diversity RL. The trained diffusion models compress the parameters of various policies into the latent space while maintaining the ability to generate policies under certain conditions. Mazoure et al. (2023) estimate value functions with diffusion models by learning the discounted state occupancy, combined with a learned reward estimator. Then, the value function can be directly computed by definition, where future states are sampled from the diffusion model. Venkatraman et al. (2023) follow Latent Diffusion Models Rombach et al. (2022) by first encoding the high-level trajectories into semantically rich representations, then applying diffusion models on them. Conditioning on latent representations, Q-functions, and policies achieves higher capability without significant extrapolation errors.
5 Applications of Diffusion Models
Diffusion models have recently been employed to address various problems in RL and several strongly related fields. We group these applications into four categories based on the task: offline RL, imitation learning, trajectory generation, and data augmentation. For each category, we provide a brief introduction to the task, followed by a detailed explanation of how existing works use diffusion models to improve performance on that task.
5.1 Offline RL
Offline RL Levine et al. (2020) aims to learn a policy from previously collected datasets without online interaction. Assuming there is a static dataset $\mathcal{D}$ collected by some (unknown) behavior policy $\pi_{\beta}$, offline RL requires the learning algorithm to derive a policy $\pi(a|s)$ that attains the most cumulative reward, which is defined in Eq. (10). The fundamental challenge in offline RL is the distributional shift: while the function approximator (e.g., policy, value function) might be trained under one distribution, it will be evaluated on a different distribution, leading to poor performance of the learned policy. High-dimensional and expressive function approximation generally exacerbates this issue:
$\pi^{*}:=\arg\max_{\pi}\mathbb{E}_{\tau\sim p_{\pi}(\tau)}\Big{[}\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})\Big{]}~{}.$ (10)
Several methods use diffusion models to help tackle or avoid the above challenge. Janner et al. (2022) first propose to generate optimal trajectories through iteratively denoising with classifier-guided sampling. Subsequent works Wang et al. (2023); Chen et al. (2023a); He et al. (2023b); Ada et al. (2023); Brehmer et al. (2023); Hansen-Estruch et al. (2023); Venkatraman et al. (2023) represent the policy as a diffusion model to capture multi-modal distributions and enhance the expressiveness of the policy class, which is beneficial to relieve the approximation error between the cloned behavior policy and true behavior policy. Ajay et al. (2023) skip the risk of distribution shift by generating state sequences with conditional diffusion models followed by inverse dynamic functions to derive executable actions, which propose a novel approach to use classifier-free guidance with low-temperature sampling to denoise out return-maximizing trajectories. In order to improve the generation ability of diffusion models for RL, Lu et al. (2023a) propose a new guidance method named contrastive energy prediction and Hu et al. (2023) capture more temporal conditions. By incorporating control-theoretic invariance into the diffusion dynamics, SafeDiffuser Xiao et al. (2023) guarantees the safe generation of planning trajectories. HDMI Li et al. (2023a) leverages a hierarchical structure to tackle long-horizon decision-making problems, which uses a reward-conditional model to discover sub-goals and a goal-conditional model to generate actions. AlignDiff Dong et al. (2023) conditions on behavior attributes with classifier-free guidance to plan to match desired trajectories accurately. CPQL Chen et al. (2023d) leverages consistency models for fast training and sampling, while EDP Kang et al. (2023) obtains speed-up by using single-step model predictions as action approximations. Diffusion models can also be used to extract the reward function Nuti et al. (2023) or value function Mazoure et al. (2023), and are prominent to estimate gradients with score-matching to solve offline optimization problems Suh et al. (2023). Enabling RL agents to generalize to multi-task and multi-agent scenarios remains a challenge due to their inherent complexity. Recent research has made progress in using diffusion models to improve the performance of policies in multi-task and multi-agent offline RL.
Multitask offline RL.
Diffusion model is verified to have the potential to address the challenge of multi-task generalization in RL. He et al. (2023a) first extend the conditional diffusion model to be capable of solving multitask decision-making problems and synthesizing useful data for downstream tasks. LCD Zhang et al. (2023a) leverages a hierarchical structure to achieve long-horizon multi-task control. Ni et al. (2023) and Liang et al. (2023) extend the idea of Diffuser Janner et al. (2022) into more specific sittings. MetaDiffuser Ni et al. (2023) demonstrates that incorporating the conditional diffusion model into the context of task inference outperforms previous meta-RL methods. AdaptDiffuser Liang et al. (2023) combines bootstrapping and diffusion-based generative modeling together to enable the model to adapt to unseen tasks.
Multi-agent offline RL.
Employing diffusion models to multi-agent RL helps model discrepant behaviors among agents and reduces approximation error. MADiff Zhu et al. (2023) uses an attention-based diffusion model to model the complex coordination among behaviors of multiple agents, which is well-suited to learning complex multi-agent interactions. DOM2 Li et al. (2023c) incorporates the diffusion model into the policy classes to enhance learning and makes it possible to generalize to shifted environments well.
5.2 Imitation Learning
The goal of imitation learning (IL) is to reproduce behavior similar to experts in the environment by extracting knowledge from expert demonstrations. Recently, many works Hegde et al. (2023); Ng et al. (2023); Chen et al. (2023c); Kapelyukh et al. (2022) have demonstrated the efficacy of representing policies as diffusion models to capture multi-modal behavior. Pearce et al. (2023) apply diffusion models to imitate human behavior in sequential environments, in which diffusion models are compared with other generative models and viable approaches are developed to improve the quality of behavior sampled from diffusion models. Chi et al. (2023); Xian et al. (2023) generate the robot’s behavior via a conditional denoising diffusion process on robot action space. Experiment results show that Diffusion models are good at predicting closed-loop action sequences while guaranteeing temporal consistency Chi et al. (2023). Li et al. (2023b) improve the models in Chi et al. (2023) by incorporating an auxiliary reconstruction loss on intermediate representations of the reverse diffusion process. Beneficial from its powerful generation ability, leveraging diffusion models to acquire diverse skills to handle multiple manipulation tasks is promising Chen et al. (2023c); Mishra et al. (2023); Xu et al. (2023); Ha et al. (2023). Diffusion models are already applied to goal-conditioned RL: Reuss et al. (2023) use a decoupled score-based diffusion model to learn an expressive goal-conditional policy. In contrast, Sridhar et al. (2023) build a unified diffusion policy to solve both goal-directed navigation and goal-agnostic exploration problems.
5.3 Trajectory Generation
Trajectory generation aims to derive a dynamically feasible path that satisfies a set of constraints. In particular, we focus on generating human pose and robot interaction sequences, which are more related to the decision-making scenario. Many works Zhang et al. (2022); Jiang et al. (2023); Tevet et al. (2022); Zhang et al. (2023b); Chen et al. (2022); Dabral et al. (2022) have remarked that the conditional diffusion models perform better than traditional methods which use GAN or Transformer. Employing a denoising-diffusion-based framework, they achieve diverse and fine-grained motion generation with various conditioning contexts Chen et al. (2023b); Carvalho et al. (2023). Recent works Du et al. (2023b); Ko et al. (2023); Du et al. (2023a) harness diffusion models to synthesize a set of future frames depicting its planned actions in the future, after which control actions are extracted from the generated video. This approach makes it possible to train policies solely based on RGB videos and deploy learned policies to various robotic tasks Black et al. (2023); Gao et al. (2023). UniSim Yang et al. (2023) uses diffusion models to build a universal simulator of real-world interaction by learning through combined diverse datasets. It can be utilized to train both high-level vision-language planners and low-level RL policies, demonstrating powerful emulation ability.
5.4 Data Augmentation
Diffusion models have already been verified to be useful for data augmentation in the RL domain. Since diffusion models perform well in learning over multimodal or even noisy distributions, they can model original data distribution precisely. What is more, they are capable of generating diverse data points to expand original distribution while maintaining dynamic accuracy. Recent works Yu et al. (2023); Chen et al. (2023e) consider augmenting the observations of robotic control using a text-guided diffusion model while maintaining the same action. The recently proposed SynthER Lu et al. (2023b) and MTDiff-s He et al. (2023a) generate complete transitions of trained tasks via a diffusion model. It proves that such augmentation brings about significant policy improvement for both online and offline RL.
6 Limitations and Remarks
In this section, we list three limitations when applying diffusion models in RL, and include remarks on comparing diffusion-based generative modeling in RL to transformer-based autoregressive approaches.
Application in online RL.
Though diffusion models have contributed to offline RL from various perspectives, the dynamic and evolving nature of online RL introduces a significant challenge, as the data distribution sampled by the current policy can change over time. The primary concern is that effectively adapting diffusion models to changing data distributions requires a substantial amount of new data Lu et al. (2023b). These models are typically trained on fixed datasets, and incorporating enough data to ensure they can generalize and remain effective in dynamic real-world scenarios is a resource-intensive task. Balancing the need for adaptability with the requirement for extensive data is the primary consideration when applying diffusion models in online RL. It is promising to solve this dilemma with more lightweight diffusion models which can keep consistency as the data distribution changes during online interactions.
Iterative sampling cost.
The unique stepwise denoising mechanism of diffusion models makes the sampling procedure need to infer the trained model multiple times, significantly increasing the sampling cost. Although we can perform sampling acceleration techniques such as DDIM or DPM-Solver, the expensive inference cost limits the model to conduct high-frequency output in online interactions with virtual environments and real-world continuous control scenarios. The problem becomes more severe in those methods that generate a long trajectory but only take the first state or action for execution, such as Diffuser Janner et al. (2022) or Decision Diffuser Ajay et al. (2023). Chen et al. (2023d) incorporate the recently proposed Consistency Model Song et al. (2023) to enable sampling with one or two diffusion steps, and the performances are comparable to that of DDPM or DDIM with 50 steps.
Variance in stochastic sampling.
In traditional RL algorithms, a continuous control policy is represented by a state-conditioned Gaussian Haarnoja et al. (2018); Schulman et al. (2017). We can take its mean as action when deterministic execution is required. However, such deterministic policy is not possible when using diffusion models as the policy class. The randomness of diffusion sampling comes from both initial noise and per-step stochastic denoising. Although per-step stochasticity can be avoided if sampling with ODE-based methods like DDIM, the randomness in initial noisy samples remains inevitable. The high-variance policies can have a negative impact in environments with high accuracy or safety requirements. Existing works in RL rarely discuss this limitation, and sampling methods with reduced variances are expected.
Comparison to transformer-based methods.
The idea of using diffusion models commonly abstracts offline RL as a conditional generative modeling problem, differing from traditional RL approaches, which need value function approximations or policy gradient calculation. This supervised learning paradigm is similar to Decision Transformer (DT) Chen et al. (2021), a sequential modeling framework for RL tasks with transformer architecture. The essence of both approaches is to take advantage of the high expressiveness model, where diffusion models enjoy a strong distribution fitting ability to produce multi-modal, diverse, and accurate outputs, and transformers are adept in long-horizon sequence modeling and time correlation understanding. Qualitatively, this difference in expertise makes the diffusion-based approach more suitable for learning complex multi-modal tasks, and Transformer-based approaches Meng et al. (2021); Wen et al. (2022) are more preferred in correlated sequence modeling in time or agent (in multi-agent tasks) dimension.
7 Summary and Future Prospects
This survey offers a comprehensive overview of contemporary research endeavors concerning the application of diffusion models in the realm of RL. According to the roles played by diffusion models, we categorize existing methodologies into using diffusion models as planners, policies, data synthesizers, and less popular roles such as value functions, latent representation models, etc. By comparing each class of methods to traditional solutions, we can see how the diffusion model addresses some of the longstanding challenges in RL, i.e., restricted expressiveness, data scarcity, compounding error, and multitask generalization. Notwithstanding these merits, it is imperative to acknowledge the existence of non-negligible limitations in using diffusion models in RL due to some inherent properties in the training and sampling of diffusion models. It is worth emphasizing that the incorporation of diffusion models into RL remains an emerging field, and there are many research topics worth exploring. Here, we outline three prospective research directions, namely, retrieval-enhanced generation, integrating safety constraints, and composing different skills.
Retrieval-enhanced generation.
Retrieval techniques are employed in various domains such as recommender systems Qin et al. (2020) and large language models Kandpal et al. (2023) to enhance the model capacity and handle long-tail distributed datasets. Some works utilize retrieved data to boost text-to-image and text-to-motion diffusion generation Sheynin et al. (2022); Zhang et al. (2023b), promoting better coverage of uncommon condition signals. During online interactions, RL agents may also encounter states that are rare in the training dataset. By retrieving relevant states as model inputs, the performance of diffusion-based decision models can be improved in these states. Also, if the retrieval dataset is constantly updated, diffusion models have the potential to generate new behaviors without retraining.
Integrating safety constraints.
Utilizing RL models for practical applications often necessitates compliance with various safety constraints García and Fernández (2015); Gu et al. (2022). Several safe RL methods transform a constrained RL problem to its unconstrained equivalent Achiam et al. (2017); Tessler et al. (2018); Sootla et al. (2022), which is then solved by generic RL algorithms. Policies acquired through these methods remain tailored to the specific constraint threshold utilized during training. Recent research Liu et al. (2023); Zhang et al. (2023c) has extended the applicability of decision transformers to the context of safety-constrained settings, thereby enabling a single model to adapt to diverse thresholds by adjusting the input cost-to-go. Similarly, diffusion models exhibit substantial potential for deployment in the domain of safe RL. Ajay et al. (2023) demonstrate that a diffusion-based planner can combine different movement skills to produce new behaviors. In addition, classifier-guided sampling can add new conditions to generated samples simply by learning additional classifiers, while the parameters of the diffusion model itself remain unchanged Dhariwal and Nichol (2021). This feature makes the diffusion model promising to be effective in scenarios where new safety constraints can be included after model training.
Composing different skills.
Most current works deploy the generation ability of diffusion models on the raw state and action spaces. From the perspective of skill-based RL Shi et al. (2022); Nam et al. (2022), it is promising to break down complex tasks into smaller, more manageable sub-skills. Diffusion models excel in modeling multi-modal distributions, and since multiple sub-skills can be viewed as distinct modes within the distribution of possible behaviors, they offer a natural fit for this task. Combining with classifier guidance or classifier-free guidance, diffusion models are possible to generate proper skills to complete the facing task. Experiments in offline RL also suggest that diffusion models can share knowledge across skills and combine them up Ajay et al. (2023); He et al. (2023a), thus having the potential for zero-shot adaptation or continuous RL by composing different skills.
Acknowledgments
The work is partially supported by National Key R&D Program of China (2022ZD0114804) and National Natural Science Foundation of China (62076161). We thank Minghuan Liu, Xihuai Wang, Jingxiao Chen and Mingcheng Chen for valuable suggestions and discussions.

### 23 Feb 2024 v4

@@
 Abstract
-Diffusion models have emerged as a prominent class of generative models, surpassing previous methods regarding sample quality and training stability.
-Recent works have shown the advantages of diffusion models in improving reinforcement learning (RL) solutions, including as trajectory planners, expressive policy classes, data synthesizers, etc.
-This survey aims to provide an overview of the advancements in this emerging field and hopes to inspire new avenues of research.
-First, we examine several challenges encountered by current RL algorithms.
-Then, we present a taxonomy of existing methods based on the roles played by diffusion models in RL and explore how the existing challenges are addressed.
-We further outline successful applications of diffusion models in various RL-related tasks while discussing the limitations of current approaches.
-Finally, we conclude the survey and offer insights into future research directions, focusing on enhancing model performance and applying diffusion models to broader tasks.
-We are actively maintaining a GitHub repository for papers and other related resources in applying diffusion models in RL  [^1] https://github.com/apexrl/Diff4RLSurvey.
+Diffusion models surpass previous generative models in sample quality and training stability.
+Recent works have shown the advantages of diffusion models in improving reinforcement learning (RL) solutions.
+This survey aims to provide an overview of this emerging field and hopes to inspire new avenues of research.
+First, we examine several challenges encountered by RL algorithms.
+Then, we present a taxonomy of existing methods based on the roles of diffusion models in RL and explore how the preceding challenges are addressed.
+We further outline successful applications of diffusion models in various RL-related tasks.
+Finally, we conclude the survey and offer insights into future research directions.
+We are actively maintaining a GitHub repository for papers and other related resources in utilizing diffusion models in RL  [^2] https://github.com/apexrl/Diff4RLSurvey.
 1 Introduction
@@
 Section 5 discusses the contribution of diffusion models on different RL-related applications.
-In Section 6, we point out the limitations when applying diffusion models and compare them with the transformer-based methods.
 Section 6 summarizes the survey with a discussion on emerging new topics.
-2 Challenges in Reinforcement Learning
 Figure 1: An illustration of how diffusion models play a different role in the classic Agent-Environment-Buffer cycle compared to previous solutions.
@@
 (3) When used as a data synthesizer, diffusion models augment the dataset with generated data sampled from the learned dataset distribution, whereas augmentation with random perturbations might generate samples that deviate from data samples.
+2 Challenges in Reinforcement Learning
 In this section, we list four challenges in RL algorithms and briefly discuss why diffusion models can address them.
 2.1 Restricted Expressiveness in Offline Learning
-Online RL Sutton and Barto (2018); Arulkumaran et al.
-(2017) has been criticized for low sample efficiency, which makes it difficult to be applied in real-world scenarios.
+Due to the low-sample efficiency, online RL Sutton and Barto (2018) is challenging to be applied in real-world scenarios.
 Offline RL Fujimoto et al.
 (2019); Kumar et al.
-(2020); Fujimoto and Gu (2021), which learns optimal policies purely from pre-collected datasets, obviates the need to interact with the environment during training, and can significantly improve sample efficiency.
-Directly applying off-policy RL methods Lillicrap et al.
-(2015); Haarnoja et al.
-(2018) to offline learning suffers from the extrapolation error problem Fujimoto et al.
+(2020) learns optimal policies from pre-collected datasets without environmental interaction, which can significantly improve sample efficiency.
+Directly applying off-policy RL to offline learning causes severe extrapolation errors Fujimoto et al.
 (2019).
-Existing works either penalize the value predictions on out-of-distribution samples Kumar et al.
-(2020); Nachum et al.
-(2019), or constrain the learning policy to be close to the data collection policy Wu et al.
-(2019); Kostrikov et al.
+Existing works penalize value predictions on out-of-distribution samples Kumar et al.
+(2020) or limit the learning policy to be close to the data collecting policy Kostrikov et al.
 (2021).
-However, penalties on the value function can make the learned policy over-conservative Lyu et al.
-(2022); for algorithms using policy constraints, since the policy is usually parameterized as a unimodal Gaussian, the restricted expressiveness makes it hard to fit the possibly diversified dataset.
-The reinforcement learning via supervised learning framework (RvS) Schmidhuber (2019); Srivastava et al.
-(2019) is now becoming another essential paradigm in offline RL, which bypasses the need for Q learning and is thus free of extrapolation errors.
-RvS learns a policy conditioned on the observed returns via supervised learning and then conditions the learned policy on a high enough return during online evaluations to generate desired behaviors Chen et al.
-(2021); Lee et al.
+However, penalties imposed on the value function may result in an overly conservative policy Lyu et al.
+(2022); when imposing policy constraints on commonly used unimodal Gaussian parameterization, the restricted expressiveness makes it difficult to fit the possibly diverse dataset.
+Reinforcement learning via supervised learning framework (RvS) Schmidhuber (2019) is another important paradigm in offline RL, which eliminates Q-learning thus free of extrapolation errors.
+RvS learns a policy conditioned on the observed returns via supervised learning and then conditions it on a high return to generate desired behaviors Chen et al.
 (2021).
-Similar to policy constraining methods, RvS requires fitting the entire offline dataset.
-Therefore, the expressiveness of parametrized policies also matters in RvS.
-Since diffusion models can represent arbitrary normalizable distributions Neal and others (2011), they hold potential to effectively improve the performance of policy constraining and RvS algorithms on complex datasets.
+Similar to policy constraining, RvS requires fitting the entire dataset.
+Therefore, the expressiveness of parameterized policies also matters in RvS.
+Diffusion models have the capability to represent any normalizable distribution Neal and others (2011), with the potential to effectively improve the performance of policy constraining and RvS algorithms on complex datasets.
 2.2 Data Scarcity in Experience Replay
 Off-policy and offline RL methods use different levels of experience replay to improve sample efficiency.
-Note that experience replay in some cases only refers to data reuse in off-policy RL.
-Here, we use the term to broadly refer to updating the current model with rollout data from other policies.
-In off-policy RL, although all previously collected experiences can be used for policy learning, the available data during training may still be inadequate for effective policy optimization due to the speed limit of simulations and the potentially large state and action spaces.
-In offline RL, as no further interactions are allowed, policy learning is more limited by the quality and coverage of the dataset.
-Inspired by the success of data augmentation in computer vision, some works implement similar augmentation techniques in RL to mitigate the data scarcity problem.
+Note that experience replay in some literature only refers to data reuse in off-policy RL.
+Here, the term broadly refers to updating the current model with rollout data from other policies.
+Although all previous experiences can be used for policy learning in off-policy RL, the limitation of simulation speed and the potentially huge state and action spaces may still hinder policy optimization.
+In offline RL, policy learning is more limited by the quality and coverage of the dataset as no further interactions are allowed.
+Inspired by the success of data augmentation in computer vision, some works adopt similar augmentation in RL to reduce data scarcity.
 RAD Laskin et al.
-(2020) uses typical image augmentation techniques such as random cropping or rotation to boost the learning efficiency in vision-based RL.
+(2020) uses image augmentation such as random cropping or rotation to improve learning efficiency in vision-based RL.
 Imre (2021) and Cho et al.
 (2022) use generative models, VAE Kingma and Welling (2013) and GAN Goodfellow et al.
-(2014), to augment the real dataset with synthetic data sampled from the learned data distribution.
-However, existing works either suffer from a lack of fidelity when using random augmentation or are limited to simple environments due to insufficient modeling ability of particular generative models, making it difficult for them to extend to more complex tasks.
-Diffusion models have already demonstrated notable performances surpassing previous generative models in domains including image and video synthesis Ho et al.
-(2020, 2022).
-When applied to RL data, diffusion models are more suitable for augmenting high-dimensional datasets with intricate interactions.
+(2014), to augment the real dataset with generated synthetic data.
+However, existing works either lack fidelity when using random augmentation or are limited to simple environments due to insufficient modeling ability of particular generative models, making them difficult to be applied to more complex tasks.
+Diffusion models have demonstrated notable performances surpassing previous generative models in high-resolution image synthesis Ho et al.
+(2020).
+When applied to RL data, diffusion models are better suited for enhancing complex interactions.
 2.3 Compounding Error in Model-based Planning
-Model-based RL (MBRL) Luo et al.
-(2022); Moerland et al.
-(2023) fits a model of dynamic transitions either from data obtained from online rollout or an offline dataset, and expects the model to facilitate decision-making.
+MBRL Luo et al.
+(2022) fits a dynamic model from online rollout data or offline datasets to facilitate decision-making.
 Common dynamic models mimic single-step state transitions and rewards in the dataset.
-Due to the limited data support and the possible stochasticity of the ground-truth transition function, there could be single-step errors when predicting with a neural network model.
-As a result, the compounding error problem arises when using the model for multistep planning Xiao et al.
-(2019), i.e., cumulative single-step errors can cause the planned states to deviate from the dataset distribution, further increasing the error in the subsequent model predictions.
-In contrast, diffusion models with powerful modeling ability of joint distributions can operate on the trajectory level and plan for all time steps simultaneously, offering better temporal consistency and less compounding errors.
+When predicting with a neural dynamic model, there could be single-step errors due to the limited data support and stochastic environment transitions.
+Cumulative single-step errors can make planned states deviate from the dataset distribution, which causes the compounding error problem when using the model for multistep planning Xiao et al.
+(2019).
+In contrast, diffusion models with powerful modeling ability of joint distributions can operate on the trajectory level and plan for multiple time steps simultaneously, improving temporal consistency and reducing compounding errors.
 2.4 Generalization in Multitask Learning
-Normal RL algorithms lack generalization abilities at the task level Vithayathil Varghese and Mahmoud (2020).
+Normal RL algorithms lack generalization abilities at the task level Beck et al.
+(2023).
 Even in the same environment, changing the reward function requires retraining a policy from scratch.
-Existing works studying online multitask RL Yu et al.
-(2020); Liu et al.
-(2021) attempt to learn the same policy on different task environments, suffering from the problem of conflicting gradients across multiple tasks, as well as low sample efficiency due to pure online learning.
-Recently, it has become a trending research direction to utilize a high-capacity model trained on multitask offline datasets and then deployed on new tasks with or without online fine-tuning Taiga et al.
-(2022); Oh et al.
+Existing online multitask RL Liu et al.
+(2021) works attempt to learn the same policy in different task environments, suffering from conflicting gradients across multiple tasks and low sample efficiency due to pure online learning.
+Recently, it has been a popular research direction to train a high-capacity model on multitask offline datasets and then deploy it on new tasks with or without online fine-tuning Taiga et al.
 (2022).
 Transformer-based pre-training decision models like Gato Reed et al.
-(2022) achieve notable successes in multitask policy learning.
-However, they typically require high-quality datasets and come with large parameter sizes as well as corresponding high training and inference costs.
-How to design an algorithm that can efficiently fit multitask datasets of mixed quality and generalize to new tasks emerges as a vital issue in multitask RL.
-As a powerful class of generative models, diffusion models can handle multimodal distributions in multitask datasets, and adapt to new tasks by estimating the task distribution.
+(2022) excel at multitask policy learning.
+However, they typically require high-quality datasets, large parameter sizes, and high training and inference costs.
+In multitask RL, designing an algorithm that can fit mixed-quality multitask datasets and generalize to new tasks emerges as a vital issue.
+As a powerful class of generative models, diffusion models can deal with multimodal distributions in multitask datasets, and adapt to new tasks by estimating the task distribution.
 3 Foundations of Diffusion Models
-This section provides the foundations of diffusion models.
-Two prominent formulations are presented: Denoising Diffusion Probabilistic Model (DDPM) Ho et al.
-(2020) and Score-based Generative Models Song et al.
-(2021).
-DDPM is widely used due to its simplicity, whereas the score-based formulation extends it to encompass continuous-time diffusion processes.
-Furthermore, guided sampling methods play a critical role in integrating diffusion models into RL frameworks.
-These methods can be divided into two main categories based on their approach to guiding the sampling process: classifier-guidance Dhariwal and Nichol (2021), which requires an extra classifier, and classifier-free guidance Ho and Salimans (2022), which makes guiding conditions as part of the model input.
-Additionally, in order to boost sampling speed especially during online interactions, fast sampling techniques have been adopted when using diffusion models in RL-related tasks Kang et al.
-(2023); Chi et al.
-(2023).
-Here we briefly cover some representative works in studying fasting sampling of diffusion models, including learning-based and learning-free methods.
+We introduce the foundations of diffusion models, including two prominent formulations and several sampling techniques that are particularly important in RL tasks.
 3.1 Denoising Diffusion Probabilistic Model
-Assuming that the real data, denoted as $x^{0}$, is sampled from an underlying distribution $q(x^{0})$, the Denoising Diffusion Probabilistic Model (DDPM) utilizes a parameterized diffusion process, represented as $p_{\theta}(x^{0})=\int p(x^{T})\prod_{t=1}^{T}p_{\theta}(x^{t-1}|x^{t})\mathop{}\!\mathrm{d}x^{1:T}$, to model how the pure noise, denoted as $x^{T}=\mathcal{N}(0,I)$, is denoised into real data $x^{0}$.
-In this formulation, each step of the diffusion process is represented by $x^{t}$, with $T$ indicating the total number of steps.
-It is important to note that both diffusion models and RL use time step notations; thus, following common practice in RL, we denote diffusion steps with superscripts and RL time steps with subscripts.
-The sequence $x^{T:0}$ is defined as a Markov chain with learned Gaussian transitions characterized by
-$p_{\theta}(x^{t-1}|x^{t})=\mathcal{N}(\mu_{\theta}(x^{t},t),\Sigma(x^{t},t))~{}.$
-If the process is reversed to $x^{0:T}$, each step is the forward transition $q(x^{t}|x^{t-1})$.
-This transition can be interpreted as adding Gaussian noise to the data according to a variance schedule $\beta^{1:T}$:
-$x^{t}=\sqrt{\alpha^{t}}x^{t-1}+\sqrt{1-\alpha^{t}}\epsilon_{t}~{},$ (1)
-where $\alpha^{t}=1-\beta^{t}$, $\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
-$q$ also holds the Markov property as $q(x^{t}|x^{t-1})$ only depends on $x^{t-1}$.
-In general, $\alpha^{T}$ should satisfy $\lim_{T\rightarrow+\infty}\alpha^{T}\rightarrow 0$ to ensure that $x^{T}$ would be a pure standard Gaussian noise when $T$ is sufficiently large.
-And from Eq.
-(1), we can infer that
-$x^{t}=\sqrt{\bar{\alpha}^{t}}x^{0}+\sqrt{1-\bar{\alpha}^{t}}{\epsilon(x^{t},t)}~{},$ (2)
-where $\bar{\alpha}^{t}=\prod_{1}^{t}\alpha^{i}$ and ${\epsilon(x^{t},t)}\sim\mathcal{N}(0,\mathbf{I})$ is the integrated noise from step 0 to $t$.
-$\epsilon(x,t)$ is unknown and to be learned by a network.
-So far, we can calculate the distribution from step 0 to step $t$ by simply one calculation.
-As for the denoising process $q(x^{t-1}|x^{t})$, it is complex while $q(x^{t-1}|x^{t},x^{0})$ is relatively simpler.
-Based on Bayes Theorem and the Markov property of $q$, we have
-$q(x^{t-1}|x^{0},x^{t})=\frac{q(x^{t-1}|x^{0})}{q(x^{t}|x^{0})}q(x^{t}|x^{t-1})=\mathcal{N}(\mu,\sigma^{2})~{},$
-where
-$\sigma^{2}(x^{t},t)$ $=\beta^{t}\frac{1-\bar{\alpha}^{t-1}}{1-\bar{\alpha}^{t}},$ (3)
-$\mu(x^{t},t)$ $=\frac{1-\bar{\alpha}^{t-1}}{1-\bar{\alpha}^{t}}\sqrt{\alpha^{t}}x^{t}+\frac{\sqrt{\bar{\alpha}^{t-1}}\beta^{t}}{1-\bar{\alpha}^{t}}x^{0}~{}.$
-From Eq. (2), we can eliminate $x^{0}$, yielding:
-$\mu(x^{t},t)=\frac{1}{\sqrt{{\alpha}^{t}}}(x^{t}-\frac{\beta^{t}}{\sqrt{1-\bar{\alpha}^{t}}}\epsilon(x^{t},t))~{}.$ (4)
-This allows us to sample $x^{T}$ from standard Gaussian noise and progressively denoise it step by step until we obtain $x^{0}$.
-However, the noise variable $\epsilon$ is still unknown.
-To address this, a network $\epsilon_{\theta}$, parameterized by $\theta$, is employed to learn the noise generation process.
-In general, the loss function is the distance between the generated $\mu(x^{t},t)$ using random variables $\epsilon$ and the network output $\epsilon_{\theta}$.
-Specially, $\mu(x^{t},t)$ is defined as in Eq.
-(4), and $\mu_{\theta}(x^{t},t)=\frac{1}{\sqrt{{\alpha}^{t}}}(x^{t}-\frac{\beta^{t}}{\sqrt{1-\bar{\alpha}^{t}}}\epsilon_{\theta}(x^{t},t))$.
-Ho et al. (2020) discover that the following version of the loss function ignoring the weights has a better performance in experiments:
-$\mathcal{L}_{t-1}\coloneqq\mathbb{E}_{x^{0},\epsilon}[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}^{t}}x^{0}+(\sqrt{1-\bar{\alpha}^{t}})\epsilon,t)\|^{2}]~{}.$ (5)
-Here, $\epsilon$ is sampled from the standard Gaussian distribution $\mathcal{N}(0,\mathbf{I})$ and $\mathcal{L}_{t-1}$ represents the loss function at diffusion step $t-1$, when $t>1$.
+Assuming that the real data $x^{0}$ are sampled from an underlying distribution $q(x^{0})$, DDPM Ho et al.
+(2020) utilizes a parameterized diffusion process, represented as $p_{\theta}(x^{0})=\int p(x^{T})\prod_{t=1}^{T}p_{\theta}(x^{t-1}|x^{t})\mathop{}\!\mathrm{d}x^{1:T}$, to model how the pure noise $x^{T}=\mathcal{N}(\mathbf{0},\mathbf{I})$ is denoised into real data $x^{0}$.
+Each step of the diffusion process is represented by $x^{t}$, with $T$ indicating the total number of steps.
+Note that both the diffusion process and RL involve time steps; thus, we denote diffusion steps as superscripts and RL time steps as subscripts.
+The sequence $x^{T:0}$ is defined as a Markov chain with learned Gaussian transitions characterized by $p_{\theta}(x^{t-1}|x^{t})=\mathcal{N}(\mu_{\theta}(x^{t},t),\Sigma(x^{t},t))$.
+If the process is reversed as $x^{0:T}$, each step is defined by the forward transition $q(x^{t}|x^{t-1})$, which is formulated as adding Gaussian noise to the data according to a variance schedule $\beta^{1:T}$:
+$x^{t}=\sqrt{\alpha^{t}}x^{t-1}+\sqrt{1-\alpha^{t}}\epsilon^{t}~{},$ (1)
+where $\alpha^{t}=1-\beta^{t}$, $\epsilon^{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. From Eq. (1), we can derive a direct mapping from $x^{0}$ to $x^{t}$:
+$x^{t}=\sqrt{\bar{\alpha}^{t}}x^{0}+\sqrt{1-\bar{\alpha}^{t}}{\epsilon(x^{t},t)}~{},$
+where $\bar{\alpha}^{t}=\prod_{1}^{t}\alpha^{i}$. From Bayes theorem and relation between $x^{t}$ and $x^{0}$, we have
+$q(x^{t-1}|x^{t},x^{0})=\mathcal{N}(\frac{1}{\sqrt{{\alpha}^{t}}}(x^{t}-\frac{\beta^{t}}{\sqrt{1-\bar{\alpha}^{t}}}\epsilon(x^{t},t)),\beta^{t}\mathbf{I})~{}.$ (2)
+Eq.
+(2) allows us to sample $x^{T}$ from Gaussian noise and denoise step by step until we obtain $x^{0}$.
+However, the noise $\epsilon(x^{t},t)$ is unknown.
+To address this, a parameterized network $\epsilon_{\theta}$ is employed to predict the noise.
+Ho et al.
+(2020) propose the following simplified loss function for learning $\epsilon_{\theta}$, which is a weighted version of the evidence lower bound (ELBO):
+$\mathcal{L}(\theta)=\mathbb{E}_{x^{0},\epsilon,t}[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}^{t}}x^{0}+\sqrt{1-\bar{\alpha}^{t}}\epsilon,t)\|^{2}]~{},$ (3)
+where $\epsilon$ is sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$.
 3.2 Score-based Generative Models
-DDPM illustrates its iterative sampling process in a step-by-step manner, with each step taking place at a discrete time interval.
 Song et al.
-(2021) extend DDPM to continuous-time cases, where the sequence $x^{0},x^{1},\ldots,x^{T}$ is replaced with a continual function $x^{t},t\in[0,T]$.
-Now, the forwarding process can be described as a Stochastic Differential Equation (SDE):
+(2021) extend DDPM to continuous-time diffusion processes, and the sequence $x^{0},x^{1},\ldots,x^{T}$ is replaced with a continuous function $x^{t},t\in[0,T]$.
+The forwarding process can be described as a Stochastic Differential Equation (SDE):
 $\mathop{}\!\mathrm{d}x=f(x,t)\mathop{}\!\mathrm{d}t+g(t)\mathop{}\!\mathrm{d}w~{},$
 where $f(x,t)$ and $g(t)$ are pre-defined functions, and $\mathop{}\!\mathrm{d}w$ is the Brownian motion.
-The example DDPM process in Eq.
-(1) is represented as $f(x,t)=-\frac{1}{2}\beta(t)x$, $g(t)=\sqrt{\beta(t)}~{}.$ According to Langevin Dynamics, the reverse of the forwarding process, as the sampling process, is described by a reverse-time SDE:
-$\mathop{}\!\mathrm{d}x=[f(x,t)-g^{2}(t)\nabla_{x}\log q_{t}(x)]\mathop{}\!\mathrm{d}t+g(t)\mathop{}\!\mathrm{d}\bar{w}~{},$ (6)
-where $\bar{w}$ is the reverse Brown Motion.
-The gradient term $\nabla_{x}\log q_{t}(x)$ is called the score function of the distribution $q_{t}(x)$.
-As the score function remains unknown, score matching Song et al.
-(2019) proposes that a parameterized score-based model $\epsilon_{\theta}(x,t)$ can be trained to approximate it.
-In practice, diffusion models usually match the scaled score function $-\sigma_{t}\nabla_{x}\log q_{t}(x)$.
-Thus, a time-dependent score model can be trained by minimizing the Fisher Divergence:
-$\min_{\theta}\mathbb{E}_{x^{0}\sim q_{0},x^{t}\sim q_{0t}}\left[\|\epsilon_{\theta}(x^{t},t)+\sigma_{t}\nabla_{x^{t}}\log q_{0t}(x^{t}|x^{0})\|_{2}^{2}\right]~{}.$
-Solving the SDE in Eq.
-(6) is difficult even with a trained score model.
-Alternatively, we can solve the corresponding Probability Flow ODE (PF-ODE), whose marginal distribution at any time is equal to that given by the original SDE in Eq.
-(6):
-$\mathop{}\!\mathrm{d}x=\left[f(x,t)-\frac{1}{2}g^{2}(t)\nabla_{x}\log q_{t}(x)\right]\mathop{}\!\mathrm{d}t~{}.$ (7)
-Given a trained score model, all terms in Eq.
-(7) are known thus it can be efficiently solved by various ODE solvers such as VODE Brown et al.
+According to Langevin dynamics, the reverse of the forwarding process is described by a reverse-time SDE:
+$\mathop{}\!\mathrm{d}x=[f(x,t)-g^{2}(t)\nabla_{x}\log p_{t}(x)]\mathop{}\!\mathrm{d}t+g(t)\mathop{}\!\mathrm{d}\bar{w}~{},$
+where $\bar{w}$ is the reverse Brown motion, $p_{t}(x)$ is the probability density of $x^{t}$, and $s(x)=\nabla_{x}\log p_{t}(x)$ is called the score function of $p_{t}(x)$.
+In practice, a parameterized score model $s_{\theta}$ is adopted to estimate the score function, which can be trained by minimizing
+$\mathcal{L}(\theta)=\mathbb{E}_{x^{0},t,x^{t}}\left[\|s_{\theta}(x^{t},t)-\nabla_{x^{t}}\log p(x^{t}|x^{0})\|_{2}^{2}\right]~{}.$
+3.3 Guided Sampling Methods
+Guided sampling from diffusion models considers sampling from the conditional distribution $p(x|y)$, where $y$ is the desired attribute of generated samples.
+Two main categories of guidance are classifier-guidance and classifier-free guidance.
+Figure 2: Different roles of diffusion models in RL.
+(a) Diffusion models as the planner.
+The sampling target is a part of trajectories whose components may vary from specific tasks.
+(b) Diffusion models as the policy.
+The sampling target is the action conditioned on the state, usually guided by the Q-function via policy gradient-style guidance or directly subtracting it from the training objective.
+(c) Diffusion models as the data synthesizer.
+The sampling target is also the trajectory, and both real and synthetic data are used for downstream policy improvement.
+For better visualizations, we omit the arrows for $N$ denoising iterations in (c) and only show generated synthetic data from randomly sampled noise.
+Note that there are other roles that are less explored, and we introduce them in Section 4.4.
+Classifier guidance.
+Classifier guided sampling relies on a differentiable classifier model $p_{\phi}(y|x)$.
+Specifically, since the guidance needs to be performed on each denoising step, the classifier model $p(y|x^{t})$ is trained on noisy samples of $x$ and corresponding attribute $y$.
+The conditional reverse process can be written as
+$p_{\theta,\phi}(x^{t-1}|x^{t},y)=Zp_{\theta}(x^{t-1}|x^{t})p_{\phi}(y|x^{t-1})~{},$ (4)
+where $Z$ is the normalization factor. Dhariwal and Nichol (2021) approximate Eq. (4) by another Gaussian distribution:
+$p_{\theta,\phi}(x^{t-1}|x^{t},y)=\mathcal{N}(\mu^{t}+w\Sigma^{t}g,\Sigma^{t})~{},$ (5)
+where $g=\nabla_{x^{t}}\log p_{\phi}(y|x^{t})|_{x^{t}=\mu^{t}}$ and $w$ is the guidance scale to control the strength of conditions.
+$\mu^{t}$ and $\Sigma^{t}$ are the mean and the covariance matrix in Eq.
+(2), respectively.
+Classifier-free guidance.
+Classifier-free sampling relies on an extra conditional noise model $\epsilon_{\theta}(x^{t},y,t)$.
+In practice, the conditional and unconditional models share the same set of parameters, and the unconditional model is represented by setting $y$ as a dummy value $\varnothing$.
+Ho and Salimans (2022) state that the noise learning target in Eq.
+(3) is a scaled score function of $p(x^{t})$, i.e., $\epsilon(x^{t},t)=-{\sigma^{t}}\nabla_{x^{t}}\log p(x^{t})$ and $\sigma^{t}=\sqrt{\beta_{t}}$.
+By using Bayes theorem, we have
+$\nabla_{x^{t}}\log p(y|x^{t})=-1/\sigma^{t}(\epsilon(x^{t},y,t)-\epsilon(x^{t},t))~{}.$
+According to Eq.
+(5), we can derive the guided noise predictor as $\bar{\epsilon}_{\theta}(x^{t},y,t)=\epsilon_{\theta}(x^{t},t)-w\sigma^{t}\nabla_{x^{t}}\log p(y|x^{t})$.
+Replacing the score function with the noise model predictions, the noise used in classifier-guided sampling can be written as
+$\hat{\epsilon}_{w}(x^{t},y,t)=w\epsilon_{\theta}(x^{t},y,t)+(1-w)\epsilon_{\theta}(x^{t},t)~{}.$
+3.4 Fast Sampling Methods
+Various fast sampling methods are proposed to overcome the prolonged iterative sampling time of diffusion models.
+We summarize these methods into two categories: those that do not involve learning and those that require extra learning, and describe representative works in each category.
+Learning-free methods.
+DDIM Song et al.
+(2022) is one of the seminal works on sampling acceleration.
+It extends DDPM to a non-Markovian formulation by learning another Markov chain $q_{\theta}(x^{t-1}|x^{t},x^{0})$.
+Some high-order solvers are proposed for diffusion sampling, such as DPM-solver Lu et al.
+(2022), which provides an excellent trade-off between sample quality and sampling speed.
+With DDIM as its first-order version, DPM-solver boosts the efficiency of solving PF-ODE, outperforming common numerical ODE solvers.
+Learning-based methods.
+Learning-based sampling methods require extra training to obtain a higher sampling efficiency at a slight expense of sampling quality.
+A recent work, Truncate Diffusion Probabilistic Model (TDPM) Zheng et al.
+(2023), demonstrates that both the noising and denoising process can be early terminated to reduce the iterative steps.
+Moreover, Watson et al.
+(2021) propose a strategy to select the best $K$ time steps to maximize the training objective for the DDPMs, which also decreases the denoising steps.
+4 The Roles of Diffusion Models in RL
+Fig.
+1 illustrates how diffusion models play a different role in RL compared to previous solutions.
+Current works applying diffusion models on RL mainly fall into three categories: as planners, as policies, and as data synthesizers.
+It is essential to note that we include methods that generate action-only sequences as planners, even though some of the representative works have “policy” in their names, e.g., Diffusion Policy Chi et al.
 (2023).
-In this case, a black-box ODE solver Dormand and Prince (1980) is highly recommended because it not only produces high-quality samples but also allows us to explicitly trade-off accuracy for efficiency.
-3.3 Guided Sampling Methods
-Diffusion models with guided sampling methods care about the conditioned data distribution $p(x|y)$, which makes it possible to generate samples with attributes of the label $y$.
-Based on whether an extra classifier model is to be trained, the methods are divided into two categories: classifier guidance and classifier-free guidance.
-An advantage of classifier guidance sampling is that the classifier and the diffusion model are trained independently.
-If you already have a diffusion model, you just need to train a classifier and integrate it with your diffusion model while sampling.
-Classifier-free guidance retrains the model totally, but it is more expressive and performs better compared to classifier guidance sampling methods.
-3.3.1 Classifier Guidance
-A simple thought is to train a differentiable discriminative model that presents $p(y|x)$.
-In other words, all we need to do is to train an extra classifier $p(y|x^{t})$ trained on noisy samples $x^{t}$.
-Assume the classifier has been pre-trained.
-Then, the reversing process is described as
-$p_{\theta,\phi}(x^{t}|x^{t+1},y)=Zp_{\theta}(x^{t}|x^{t+1})p_{\phi}(y|x^{t})~{},$ (8)
-where $Z$ is a normalization factor. Dhariwal and Nichol (2021) state that Eq. (8) can be approximately regarded as another Gaussian distribution:
-$p(x^{t}|x^{t+1},y)=\mathcal{N}(\mu(x^{t},t)+s\Sigma(x^{t},t)g,\Sigma(x^{t},t))~{},$
-where $g=\nabla_{x^{t}}\log p_{\phi}(y|x^{t})|_{x^{t}=\mu}$ and $s$ is the guidance scale to control the effect of conditions.
-The $\mu(x^{t},t)$ and $\Sigma(x^{t},t)$ function is identical to those in DDPM (Eq.
-(4) and Eq.
-(3)).
-Therefore, the sampling process can be explicitly presented as follows:
-$x^{t+1}\sim\mathcal{N}(\mu+s\Sigma g,\Sigma)~{},$
-where $\Sigma$ is short for $\Sigma(t,x^{t})$.
-3.3.2 Classifier-free Guidance
-Rather than estimating the conditioned data distribution $p(x|y)$ directly, classifier-free methods try to predict the score function $\nabla_{x}\log p(x|y)$.
-With the Bayes Theorem, we can decompose the score function into an unconditional term and a classifier conditioning term as
-$\nabla_{x}\log p(x|y)=\nabla_{x}\log p(y|x)+\nabla_{x}\log p(x)~{}.$ (9)
-Unlike the classifier guidance, the original training setup is modified so the diffusion model should be retrained.
-The network to estimate noise is defined as $\epsilon(x^{t},c)$ ($c=\varnothing$ for unconditional inputs).
-In practice, the conditional and unconditional models are trained with the same set of network parameters by randomly setting $c=\varnothing$ with a pre-specified probability during training.
-From Eq.
-(9) we can infer
-$\nabla_{x}\log p(y|x)=\nabla_{x}\log p(x|y)-\nabla_{x}\log p(x)~{}.$
-As Song et al.
-(2022) reveal, diffusion models and the score functions are equivalent, which indicates $\nabla_{x^{t}}\log p(x^{t})\propto\epsilon(x^{t},t)$.
-As a result, we can now substitute $\nabla_{x}\log p(y|x)$ with this into the formula for classifier guidance
-$\bar{\epsilon}_{w}(x^{t},y)$ $=\epsilon_{\theta}(x^{t},y)+w(\epsilon_{\theta}(x^{t},y)-\epsilon_{\theta}(x^{t}))$
-$=(1+w)\epsilon_{\theta}(x^{t},y)-w\epsilon_{\theta}(x^{t})~{},$
-where $w$ stands for the guidance scale.
-3.4 Fast Sampling Methods
-Diffusion model has been criticized for its prolonged iterative sampling time.
-Several fast sampling skills are proposed to solve this issue.
-Generally, the skills extend the diffusion model to a more fundamental paradigm to acquire efficiency.
-We categorize the methods into two categories: those that do not involve learning (learning-free) and those that require extra learning sessions (learning-based).
-Learning-free sampling methods.
-Denoising Diffusion Implicit Models (DDIM) Song et al.
-(2022) is one of the earliest works on sampling acceleration.
-It aims at extending DDPM to a non-Markovian case by learning another Markov chain $q_{\theta}(x^{t-1}|x^{t},x^{0})$.
-The work also reveals that DDPM and DDIM are special cases of a more general paradigm.
-Moreover, it is observed in later works that DDIM is the discrete version of solving the PF-ODE Eq.
+Generating multi-step action sequences can be viewed as planning in action space, and the use of diffusion models to ensure temporal consistency is similar to other planning-based diffusion methods.
+The following subsections will illustrate overall frameworks and representative papers for each category.
+4.1 Planner
+Planning in RL refers to using a dynamic model to make decisions imaginarily and selecting the appropriate action to maximize cumulative rewards.
+This process usually explores various sequences of actions and states, thus improving decisions over a longer horizon.
+Planning is commonly used in the MBRL framework with a learned dynamic model.
+However, the planning sequences are usually simulated autoregressively, which may lead to severe compounding errors, especially in the offline setting due to limited data support.
+Diffusion models offer a promising alternative as they can generate multi-step planning sequences simultaneously.
+A general framework of diffusion planners is shown in Fig.
+2(a).
+Diffusion models are designed to generate clips of the trajectory $\tau=(s_{1},a_{1},r_{1},\ldots,s_{H},a_{H},r_{H})$, denoted as $x(\tau)=(e_{1},e_{2},\ldots,e_{H})$.
+$H$ is the planning horizon.
+Here $e_{t}$ represents the selected elements from $(s_{t},a_{t},r_{t})$, where various choices can be made as $e_{t}=(s_{t},a_{t})$ Janner et al.
+(2022); Liang et al.
+(2023a); He et al.
+(2023a), $e_{t}=(s_{t},a_{t},r_{t})$ He et al.
+(2023a); Hu et al.
+(2023), $e_{t}=s_{t}$ Ajay et al.
+(2023); Zhu et al.
+(2023), or $e_{t}=a_{t}$ Chi et al.
+(2023); Li et al.
+(2023b).
+RL datasets often contain trajectory data of varying quality.
+In order to make the diffusion planner generate high rewarded trajectories during evaluations, guided sampling techniques are widely adopted.
+The guidance can be either injected in the sampling stage following the classifier guided sampling or in both the training and sampling stages following the classifier-free sampling.
+When deploying the trained diffusion planner for online evaluations, fast sampling methods are usually adopted to reduce the inference time.
+Besides, to ensure the planned trajectory is congruous with the agent’s current state, before each denoising step, the first $h\,(h\geq 1)$ steps of the noisy trajectory are substituted with the $h$ steps of historical states observed by the agent.
+Here, $h$ is a hyperparameter where a larger $h$ can better handle partially observable and non-Markovian settings but increases the modeling complexity.
+4.2 Policy
+Based on whether they rely on a dynamic model for making decisions, RL algorithms can be categorized into MBRL and model-free RL.
+Under such classification criteria, using diffusion models as planners is more akin to MBRL, as the generated trajectories encapsulate dynamics information.
+Another perspective is that diffusion planner can be seen as a combination of policy and dynamic model Janner et al.
 (2022).
-Then, some high-order solvers emerged, such as DPM-solver Lu et al.
-(2022), which provides an excellent trade-off between sample quality and sampling speed.
-With DDIM as its first-order version, DPM-solver boosts the efficiency of solving PF-ODE, outperforming common numerical ODE solvers like Runge-Kutta.
-Consequently, DPM-solver has become one of the most frequently used fast sampling methods.
-Learning-based sampling methods.
-Learning-based sampling is another approach to fast sampling.
-Unlike the learning-free methods, they include extra learning processes to reach a higher sampling efficiency at a slight expense of sampling quality.
-A recent work, Truncate Diffusion Probabilistic Model (TDPM) Zheng et al.
-(2023), proposes that both the diffusion and denoising process can be truncated so that the iterative steps are reduced.
-Specifically, the forward process is truncated when the sample is noisy enough, and the denoising process starts from a relatively noisy sample (not pure noise), which will be learned by a network or else.
-Moreover, Watson et al.
-(2021) learn a strategy to select the best $K$ time steps to maximize the training objective for the DDPMs, which also decreases the denoising steps.
+In contrast, using diffusion models as policies focuses on improving existing model-free RL solutions.
+Section 2.1 states the main drawbacks of current offline policy learning methods: over-conservatism and lack of expressiveness.
+Many works use diffusion models as the policy class in model-free RL to tackle these problems.
+Diffusion-QL Wang et al.
+(2023) first combines the diffusion policy with the Q learning framework and finds that it can perfectly fit on datasets collected by strong multi-modal behavior policies, where previous distance-based policy regularization approaches fail.
+Compared with using diffusion models as planners, the diffusion target of the diffusion policy is simply the action given the current state, as shown in Fig.
+2(b).
+Suppose the noise predictor is $\epsilon_{\theta}(a^{k},k,s)$ parameterized by $\theta$, and the derived action mean is $\mu_{\theta}(a|s)$.
+To guide the model sampling actions that can lead to high returns, it is necessary to take $Q(s,a)$ into consideration.
+Diffusion-QL includes a weighted Q maximization term into the diffusion training loss as
+$\mathcal{L}(\theta)$ $=\mathbb{E}_{k,\epsilon,(s,a)\sim\mathcal{D}}[\|\epsilon-\epsilon_{\theta}(a^{k},s,k)\|_{2}^{2}]$
+$\quad-\frac{\eta}{\mathbb{E}_{(s,a)\sim\mathcal{D}}[Q(s,a)]}\cdot\mathbb{E}_{s\sim\mathcal{D},a^{0}\sim\pi_{\theta}(\cdot|s)}[Q(s,a^{0})]~{},$
+where $\eta$ is a hyperparameter, and $\mathcal{D}$ is the offline dataset.
+Some works Chen et al.
+(2023a); Lu et al.
+(2023a); Hansen-Estruch et al.
+(2023); Kang et al.
+(2023) construct the policy by (advantage) weighted regression as
+$\pi_{\theta}(a|s)\propto\mu_{\theta}(a|s)\exp(\alpha Q(s,a))~{},$
+where $\alpha$ is the temperature hyperparameter.
+Following this, Chen et al.
+(2023a) decouple the policy learning into behavior learning and action evaluation, which allows more freedom in the choice of guidance.
+They also propose in-sample planning for Q-learning, avoiding extrapolation errors in previous offline RL methods.
+CEP Lu et al.
+(2023a) further extends this framework to sample from the more general energy-guided distribution $p(x)\propto q(x)\exp(-\beta\mathcal{E}(x))$.
+Here, $\mathcal{E}(x)$ is an energy function, and in the RL setting, it is trained via contrastive learning to match the in-sample softmax of Q functions.
+Since there are off-the-shelf Q functions or energy functions after training, some methods use those functions to further improve the sampling quality during evaluations.
+They first sample multiple candidate actions for a given state and use Q or energy values to perform weighted sampling or just select the best candidate.
+Fast reaction is crucial when deploying policies in online environments.
+Therefore, almost all diffusion policies use smaller diffusion steps during sampling, usually about 15 steps.
+ODE solvers such as the DPM-solver Lu et al.
+(2022) are also used to accelerate sampling Chen et al.
+(2023a); Lu et al.
+(2023a); Kang et al.
+(2023); Li et al.
+(2023c).
+Kang et al.
+(2023) introduce action approximation, which allows one-step action sampling in the training stage.
+4.3 Data Synthesizer
+In addition to fitting multi-modal distributions, a simple and common use of diffusion models is to create synthetic data, which has been widely applied in computer vision.
+Therefore, the diffusion model is a natural data synthesizer for RL datasets because data scarcity is a practical concern.
+To ensure consistency of synthetic data to the environment dynamics, previous data augmentation approaches in RL usually add minor perturbations to states and actions Sinha et al.
+(2021).
+In contrast, Fig.
+2(c) illustrates that diffusion models generate diverse and consistent data by learning the data distribution from the entire dataset $\mathcal{D}_{\mathrm{real}}$.
+The diffusion model first learns the parameterized data distribution $\rho_{\theta}(\tau)$ from the real data $\mathcal{D}_{\mathrm{real}}$, and generates desired synthetic data by
+$\mathcal{D}_{\mathrm{syn}}=\{\tau\sim\rho_{\theta}(\tau)\}~{}.$
+Then, real and synthetic data are combined together as
+$\mathcal{D}=\mathcal{D}_{\mathrm{real}}\cup\mathcal{D}_{\mathrm{syn}}~{},$
+and $\mathcal{D}$ is used for policy learning.
+In online settings, the policy interacts with the environment, collects more real data into $\mathcal{D}_{\mathrm{real}}$, and updates the diffusion model.
+As a result, the diffusion model and the policy are updated alternately.
+4.4 Others
+Besides the primary directions discussed above, other ways of improving RL with diffusion models are less explored.
+Mazoure et al.
+(2023) estimate value functions with diffusion models by learning the discounted state occupancy, combined with a learned reward estimator.
+Then, the value function can be directly computed by definition, where future states are sampled from the diffusion model.
+Venkatraman et al.
+(2023) first encode the high-level trajectories into semantically rich latent representations, then apply diffusion models to learn the latent distribution.
+Conditioning on latent representations improves the capability of Q-functions and policies without significant extrapolation errors.
+Rigter et al.
+(2023) use a diffusion dynamic model and allow an online RL policy to collect synthetic trajectories on it.
+The interplay of the diffusion dynamic model and RL policy is done by alternating between state denoising and Langevin dynamics of policy actions.
 Table 1: Summary of papers on diffusion models for RL.
@@
 Diffuser Janner et al. (2022) Planner Offline Classifier
-AdaptDiffuser Liang et al. (2023) Offline Classifier
+AdaptDiffuser Liang et al. (2023a) Offline Classifier
 EDGI Brehmer et al. (2023) Offline Classifier
 TCD Hu et al. (2023) Offline Classifier-free
-Crossway Diffusion Li et al. (2023b) Offline; Robotics None
-SGP Suh et al. (2023) Offline; Robotics None
-GSC Mishra et al. (2023) Imitation; Robotics Classifier
-UniPi Du et al. (2023a) Imitation; Robotics Classifier-free
-ChainedDiffuser Xian et al. (2023) Imitation; Robotics None
+LatentDiffuser Li (2023) Offline Energy Function
+HDMI Li et al. (2023a) Offline; Hierarchical Classifier-free
+SafeDiffuser Xiao et al. (2023) Offline; Safe None
+MADiff Zhu et al. (2023) Offline; Multi-agent Classifier-free
+MTDiff-p He et al. (2023a) Offline; Multitask Classifier-free
+MetaDiffuser Ni et al. (2023) Offline; Multitask Classifier-free
 Diffusion Policy Chi et al. (2023) Imitation; Robotics None
+Crossway Diffusion Li et al. (2023b) Imitation; Robotics None
 AVDC Ko et al. (2023) Imitation; Robotics None
-MTDiff-p He et al. (2023a) Offline; Multi-task Classifier-free
-MetaDiffuser Ni et al. (2023) Offline; Multi-task Classifier-free
-SafeDiffuser Xiao et al. (2023) Offline; Safe Self-proposed
-MADiff Zhu et al. (2023) Offline; Multi-agent Classifier-free
-HDMI Li et al. (2023a) Offline; Hierarchical Classifier-free
+SkillDiffuser Liang et al. (2023b) Imitation; Multitask; Hierarchical Classifier-free
 MLD Chen et al. (2022) Trajectory Generation Classifier-free
 MDM Tevet et al. (2022) Trajectory Generation Classifier-free
-UniSim Yang et al. (2023) Trajectory Generation Classifier-free
+UniSim Yang et al. (2023b) Trajectory Generation Classifier-free
 ReMoDiffuse Zhang et al. (2023b) Trajectory Generation Classifier-free
@@
 SRDP Ada et al. (2023) Offline Q-loss
-SfBC Chen et al. (2023a) Offline Policy Gradient
-IDQL Hansen-Estruch et al. (2023) Offline Policy Gradient
-EDP Kang et al. (2023) Offline Policy Gradient
-DiffCPS He et al. (2023b) Offline None
-NoMaD Sridhar et al. (2023) Robotics None
-BESO Reuss et al. (2023) Offline; Goal-conditioned Classifier-free
-CEP Lu et al. (2023a) Offline; Image Synthesis Policy Gradient
+EDP Kang et al. (2023) Offline Q-loss
+SfBC Chen et al. (2023a) Offline Sample & Reweight
+IDQL Hansen-Estruch et al. (2023) Offline Sample & Reweight
+DiffCPS He et al. (2023b) Offline Convex Optimization
+CPQL Chen et al. (2023d) Offline; Online Q-loss
+CEP Lu et al. (2023a) Offline; Image Synthesis Energy Function
 DOM2 Li et al. (2023c) Offline; Multi-agent Q-loss
-CPQL Chen et al. (2023d) Offline; Online Q-loss
+NoMaD Sridhar et al. (2023) Imitation; Robotics None
+BESO Reuss et al. (2023) Imitation; Goal-conditioned Classifier-free
  Pearce et al. (2023) Imitation Classifier-free
@@
 SynthER Lu et al. (2023b) Offline; Online None
-MTDiff-s He et al. (2023a) Offline; Multi-task Classifier-free
+MTDiff-s He et al. (2023a) Offline; Multitask Classifier-free
 LDCQ Venkatraman et al. (2023) Latent Representation Offline Classifier-free
-LatentDiffuser Li (2023) Offline Policy Gradient
- Hegde et al. (2023) Quality Diversity None
 DVF Mazoure et al. (2023) Value Function Offline None
-4 The Roles of Diffusion Models in RL
-Figure 2: Different roles of diffusion models in RL.
-(a) Diffusion models as the planner.
-The sampling target is a part of trajectories whose components may vary from specific tasks.
-(b) Diffusion models as the policy.
-The sampling target is the action conditioned on the state, usually guided by the Q-function via policy gradient-style guidance or directly subtracting it from the training objective.
-(c) Diffusion models as the data synthesizer.
-The sampling target is also the trajectory, and both real and synthetic data are used for downstream policy improvement.
-For better visualizations, we omit the arrows for $N$ denoising iterations in (c) and only show generated synthetic data from randomly sampled noise.
-Note that there are other roles that are less explored, and we introduce them in Section 4.4.
-Diffusion models have proven their ability to generate diverse data and model multi-modal distributions.
-Considering the long-existing challenges introduced in Section 2, it is sufficient to improve the performance and sample efficiency of RL algorithms with diffusion models.
-In Fig.
-1, we illustrate how diffusion models play a different role in RL compared to previous solutions.
-Current works applying diffusion models on RL mainly fall into three categories: using diffusion models as the planner, as the policy, and as the data synthesizer.
-It is essential to note that we include methods that generate action-only sequences as planners, even though some of the representative works have “policy” in their names, e.g., Diffusion Policy Chi et al.
-(2023).
-Generating multi-step action sequences can be viewed as planning in action space, and the use of diffusion models to ensure temporal consistency is similar to other planning-based diffusion methods.
-The following subsections will illustrate overall frameworks and representative papers for each category.
-4.1 Planner
-Planning in RL refers to the process of using a model of the environment to make decisions imaginarily, and then selecting the best action in order to maximize a cumulative reward signal.
-This process usually simulates or explores different sequences of actions and states, predicting the outcomes of its decisions, thus resulting in better actions from the perspective of a longer horizon.
-Therefore, planning is commonly applied in the MBRL framework.
-However, the decision sequences used for planning are generated autoregressively, which may lead to severe compounding errors, especially in the offline setting, due to the limited data support.
-Diffusion models provide a possible solution since they can generate the whole sequence simultaneously.
-A general framework of diffusion models as planners is shown in Fig.
-2(a), which was first proposed by Janner et al.
-(2022).
-Inputs and outputs of the diffusion model are usually clips of the real trajectory $\tau=(s_{1},a_{1},r_{1},\ldots,s_{H},a_{H},r_{H})$, denoted as $x(\tau)=(e_{1},e_{2},\ldots,e_{H})$.
-Here $e_{t}$ represents for the selected elements from $(s_{t},a_{t},r_{t})$, where various choices can be made as $e_{t}=(s_{t},a_{t})$ Janner et al.
-(2022); Liang et al.
-(2023); He et al.
-(2023a); Xiao et al.
-(2023), or $e_{t}=(s_{t},a_{t},r_{t})$ He et al.
-(2023a); Hu et al.
-(2023), $e_{t}=s_{t}$ Ajay et al.
-(2023); Zhu et al.
-(2023), or $e_{t}=a_{t}$ Chi et al.
-(2023); Li et al.
-(2023b).
-$y(\tau)$ is the guidance that contains the desired property of the generated trajectory, such as the discounted return, the indicator of whether the task is completed, etc.
-Such guidance is meant to lead the diffusion model to generate good trajectories instead of those that only satisfy the environment dynamics.
-Suppose the diffusion model $\epsilon_{\theta}$ parameterized by $\theta$ predicts the noise added in the forward process, and $x^{k}\coloneqq x(\tau^{k})$ is the corresponding clip of $\tau$ at the $k$-th diffusion step.
-Either classifier guidance or classifier-free guidance is incorporated.
-For classifier guidance Janner et al.
-(2022); Liang et al.
-(2023), $\epsilon_{\theta}$ is trained as usual by taking $\tau^{k}$ and the timestep $k$ as input:
-$\mathcal{L}(\theta)=\mathbb{E}_{\tau,k,\epsilon}[\|\epsilon-\epsilon_{\theta}(x^{k},k)\|_{2}^{2}]~{}.$
-After training, suppose $\mu_{\theta}(x^{k},k)$ and $\Sigma$ are the mean and the variance of the original reverse process $p_{\theta}(x^{k-1}|x^{k})$, respectively.
-The gradient of $y(\tau)$ is injected during sampling as the guidance:
-$p_{\theta}(x^{k-1}|x^{k},y(\tau))\approx\mathcal{N}(x^{k-1};\mu_{\theta}+\alpha\Sigma^{k}\nabla y(\tau),\Sigma^{k})~{},$
-where $\alpha$ is a hyperparameter.
-In contrast, classifier-free guidance Ajay et al.
-(2023); He et al.
-(2023a); Hu et al.
-(2023) injects $y(\tau)$ in both training and sampling stages.
-The noise prediction model $\epsilon_{\theta}$ now takes an additional $y(\tau)$ as input, giving the training objective:
-$\mathcal{L}(\theta)=\mathbb{E}_{\tau,k,\epsilon,\beta}[\|\epsilon-\epsilon_{\theta}(x^{k},(1-\beta)y(\tau)+\beta\varnothing,k)\|_{2}^{2}]~{},$
-where $\beta$ is sampled from a Bernoulli distribution with probability $p$. At sampling, $y(\tau)$ is directly fed into $\epsilon_{\theta}$, as
-$p_{\theta}(x^{k-1}|x^{k},y(\tau))\approx\mathcal{N}(x^{k-1};\mu_{\theta}(x^{k},y(\tau),k),\Sigma^{k})~{}.$
-As the time consumption of sampling using the original DDPM is unacceptable when deploying RL in the online environment, fast sampling methods are commonly incorporated, and the diffusion step for sampling is usually set to 100 or 200.
-Besides, to guarantee sample stability during the sampling process, the first several steps of a sampled trajectory can be replaced with the ground truth as a hard constraint.
-For example, when planning the trajectory after $e_{1:h}$, the model first generates the full trajectory $\tilde{x}^{k}=\tilde{e}_{1:H}^{k}$ at the $k$-the diffusion step.
-Then the first $h$ steps are replaced, forming $\tilde{x}^{k}\leftarrow(e_{1},e_{2},\ldots,e_{h},\tilde{e}_{h+1}^{k},\ldots,\tilde{e}_{H}^{k})$.
-The core ideas behind why non-autoregressive diffusion models can generate Markovian decision sequences are stated by Janner et al.
-(2022).
-With the U-Net backbone, the sampled token $e_{t}^{k}$ takes information in a local receptive field $e_{t-l:t+l}^{k+1}$, where $l$ is determined by the model structure.
-Once the local consistency is guaranteed, the global consistency will also be built during the denoising process as information spreads among the trajectory.
-Though this raises new confusion that using future information is anti-causal for sequential problems, RL includes a hidden hypothesis that the current decision will lead to optimal future outcomes.
-That is, utilizing future information does not violate the general RL framework.
-Benefiting from the non-autoregressive scheme, diffusion models surpass traditional models in generating planning sequences.
-First, such generations do not suffer from the aforementioned compounding errors.
-Then, diffusion models fit for both continuous-reward and sparse-reward settings Janner et al.
-(2022), even for more challenging multitask planning He et al.
-(2023a), safe planning Xiao et al.
-(2023), and generating based on multi-modal observations Du et al.
-(2023a); Chi et al.
-(2023).
-Besides, diffusion models can also learn from non-Markovian data by exploiting temporal information Hu et al.
-(2023).
-4.2 Policy
-Compared with traditional RL taxonomy, which roughly divides RL algorithms into MBRL and model-free RL, using diffusion models as the planner is similar to MBRL and focuses on capturing the environment dynamics.
-In contrast, taking diffusion models as the policy follows the framework of model-free RL.
-Section 2.1 states the main drawbacks of offline policy learning frameworks: over-conservatism and poor capability on diversified datasets.
-With excellent expressive ability on multi-modal distribution, many works utilize diffusion models as the policy to tackle these problems.
-Diffusion-QL Wang et al.
-(2023) first explores the advantages of the diffusion policy in offline RL and finds that it can perfectly fit on dataset generated by strong multi-modal behavior policies, where previous distance-based BC methods fail.
-Experiment results also support the dominance of diffusion policies on diversified datasets Lu et al.
-(2023a); Chen et al.
-(2023a); Hansen-Estruch et al.
-(2023); Ada et al.
-(2023); Li et al.
-(2023c).
-Compared with diffusion models as planners, the diffusion target of the diffusion policy is simply the action given the current state, as shown in Fig.
-2(b).
-Suppose the noise predictor is $\epsilon_{\theta}(a^{k},k,s)$ parameterized by $\theta$, and the derived diffusion model is $\mu_{\theta}(a|s)$.
-To guide the model sampling actions with high Q-values, it is necessary to take $Q(s,a)$ into consideration.
-Some papers Chen et al.
-(2023a); Lu et al.
-(2023a); Hansen-Estruch et al.
-(2023); Kang et al.
-(2023) construct the policy by (advantage) weighted regression as
-$\pi_{\theta}(a|s)\propto\mu_{\theta}(a|s)\exp(\alpha Q(s,a))~{},$
-where $\alpha$ is a hyperparameter.
-Following this, Chen et al.
-(2023a) decouple the policy learning into behavior learning and action evaluation, which allows more freedom in the choice of guidance.
-It also proposes in-sample planning for Q-learning, which avoids extrapolation errors in previous offline RL methods.
-Lu et al.
-(2023a) further generalize this framework to any energy reweighted distribution $p(x)\propto q(x)\exp(-\beta\mathcal{E}(x))$, where $\mathcal{E}(x)$ is an energy function, and train it via contrastive learning.
-Approaches using explicit Q-values as guidance are also considered.
-Wang et al.
-(2023) and Ada et al.
-(2023) subtract a weighted expectation of $Q(s,a)$ in the training loss as
-$\mathcal{L}(\theta)$ $=\mathbb{E}_{k,\epsilon,(s,a)\sim\mathcal{D}}[\|\epsilon-\epsilon_{\theta}(a^{k},s,k)\|_{2}^{2}]$
-$\quad-\frac{\eta}{\mathbb{E}_{(s,a)\sim\mathcal{D}}[Q(s,a)]}\cdot\mathbb{E}_{s\sim\mathcal{D},a^{0}\sim\pi_{\theta}(\cdot|s)}[Q(s,a^{0})]~{},$
-where $\eta$ is a hyperparameter, and $\mathcal{D}$ is the offline dataset.
-Fast reaction is crucial when deploying policies in online environments.
-Therefore, almost all diffusion policies use smaller diffusion steps during sampling, usually about 15 steps.
-ODE solvers such as the DPM-solver Lu et al.
-(2022) are also used to accelerate sampling Chen et al.
-(2023a); Lu et al.
-(2023a); Kang et al.
-(2023); Li et al.
-(2023c)Pearce et al.
-(2023) compare three different backbone structures of diffusion models, finding that MLP-sieve or Transformer is also sufficient in expressiveness.
-Kang et al.
-(2023) introduce action approximation, which allows one-step action sampling in the training stage.
-4.3 Data Synthesizer
-In addition to fitting multi-modal distributions, a simple and common use of diffusion models is to generate more training samples, which has been widely applied and proven in computer vision.
-Therefore, it is natural to apply the diffusion model as a data synthesizer on RL datasets, as data scarcity is a practical challenge of RL, as stated in Section 2.2.
-To guarantee consistency of synthetic data to the environment dynamics, previous data augmentation approaches in RL usually add small perturbations to existing states and actions Sinha et al.
-(2021).
-In contrast, Fig.
-2(c) illustrates that diffusion models learn the data distribution from the whole dataset $\mathcal{D}$, and enable generating highly diversified data while keeping consistency.
-Lu et al.
-(2023b) investigate the ability of diffusion models as the data synthesizer in both offline and online settings.
-It directly trains the diffusion model from the offline dataset or the online replay buffer and then generates more samples for policy improvement.
-Analysis shows that the quality of data generated by diffusion models is higher than those generated by explicit data augmentation in diversity and accuracy.
-With synthetic data, the performance of the offline policy and the sample efficiency of the online policy are significantly improved.
-He et al.
-(2023a) deploy diffusion models to augment data for multitask offline datasets and achieves better performance than those on single-task datasets.
-It claims that fitting on multiple tasks may enable implicit knowledge sharing across tasks, which also benefits from the multi-modal property of diffusion models.
-4.4 Others
-Besides directions that have been mainly focused on, some ways of improving RL with diffusion models are under exploration.
-Hegde et al.
-(2023) take a similar idea as hyper networks in meta-learning, generating parameters of policies for quality diversity RL.
-The trained diffusion models compress the parameters of various policies into the latent space while maintaining the ability to generate policies under certain conditions.
-Mazoure et al.
-(2023) estimate value functions with diffusion models by learning the discounted state occupancy, combined with a learned reward estimator.
-Then, the value function can be directly computed by definition, where future states are sampled from the diffusion model.
-Venkatraman et al.
-(2023) follow Latent Diffusion Models Rombach et al.
-(2022) by first encoding the high-level trajectories into semantically rich representations, then applying diffusion models on them.
-Conditioning on latent representations, Q-functions, and policies achieves higher capability without significant extrapolation errors.
+PolyGRAD Rigter et al. (2023) Dynamic Model Online None
 5 Applications of Diffusion Models
-Diffusion models have recently been employed to address various problems in RL and several strongly related fields.
-We group these applications into four categories based on the task: offline RL, imitation learning, trajectory generation, and data augmentation.
-For each category, we provide a brief introduction to the task, followed by a detailed explanation of how existing works use diffusion models to improve performance on that task.
+In this section, we conduct a complete review of existing works.
+We divide them into five groups based on the tasks they are applied: offline RL, online RL, imitation learning, trajectory generation, and data augmentation.
+For each group, we provide a detailed explanation of how each method uses diffusion models to handle the task.
 5.1 Offline RL
-Offline RL Levine et al.
-(2020) aims to learn a policy from previously collected datasets without online interaction.
-Assuming there is a static dataset $\mathcal{D}$ collected by some (unknown) behavior policy $\pi_{\beta}$, offline RL requires the learning algorithm to derive a policy $\pi(a|s)$ that attains the most cumulative reward, which is defined in Eq.
-(10).
-The fundamental challenge in offline RL is the distributional shift: while the function approximator (e.g., policy, value function) might be trained under one distribution, it will be evaluated on a different distribution, leading to poor performance of the learned policy.
-High-dimensional and expressive function approximation generally exacerbates this issue:
-$\pi^{*}:=\arg\max_{\pi}\mathbb{E}_{\tau\sim p_{\pi}(\tau)}\Big{[}\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})\Big{]}~{}.$ (10)
-Several methods use diffusion models to help tackle or avoid the above challenge.
+Offline RL aims to learn a policy from previously collected datasets without online interaction.
+Assuming there is a static dataset $\mathcal{D}$ collected by some (unknown) behavior policy $\pi_{\beta}$, offline RL requires the learning algorithm to derive a policy $\pi(a|s)$ that attains the most cumulative reward:
+$\pi^{*}=\arg\max_{\pi}\mathbb{E}_{\tau\sim p_{\pi}(\tau)}\Big{[}\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})\Big{]}~{}.$
+The fundamental challenge in offline RL is the distributional shift.
+This refers to the discrepancy between the dataset distribution used to train the function approximators (e.g., policies and value functions) and the distribution on which the policy is evaluated.
+This mismatch often results in subpar online performance.
+High-dimensional and expressive function approximation generally exacerbates this issue.
+Several methods use diffusion models to help tackle or avoid the above challenges.
 Janner et al.
-(2022) first propose to generate optimal trajectories through iteratively denoising with classifier-guided sampling.
+(2022) first propose to generate optimal trajectories through iterative denoising with classifier-guided sampling.
 Subsequent works Wang et al.
@@
 (2023); Hansen-Estruch et al.
-(2023); Venkatraman et al.
 (2023) represent the policy as a diffusion model to capture multi-modal distributions and enhance the expressiveness of the policy class, which is beneficial to relieve the approximation error between the cloned behavior policy and true behavior policy.
 Ajay et al.
-(2023) skip the risk of distribution shift by generating state sequences with conditional diffusion models followed by inverse dynamic functions to derive executable actions, which propose a novel approach to use classifier-free guidance with low-temperature sampling to denoise out return-maximizing trajectories.
+(2023) alleviate the distribution shift problem by generating state sequences with conditional diffusion models followed by inverse dynamic functions to derive executable actions, which propose a novel approach to use classifier-free guidance with low-temperature sampling to denoise out return-maximizing trajectories.
+LatentDiffuser Li (2023) performs diffusion planning over a learned latent space with separate decoders to recover raw trajectories.
+Benefiting from a more compact planning space, it achieves superior performances on long-horizon and high-dimensional tasks.
 In order to improve the generation ability of diffusion models for RL, Lu et al.
@@
 HDMI Li et al.
-(2023a) leverages a hierarchical structure to tackle long-horizon decision-making problems, which uses a reward-conditional model to discover sub-goals and a goal-conditional model to generate actions.
-AlignDiff Dong et al.
-(2023) conditions on behavior attributes with classifier-free guidance to plan to match desired trajectories accurately.
+(2023a) adopts a hierarchical structure to tackle long-horizon decision-making problems, which uses a reward-conditional model to discover sub-goals and a goal-conditional model to generate actions.
+Dong et al.
+(2023) condition the diffusion planner on diverse behavior attributes and learn from human preferences to generate trajectories that can match user-customized behaviors.
 CPQL Chen et al.
-(2023d) leverages consistency models for fast training and sampling, while EDP Kang et al.
-(2023) obtains speed-up by using single-step model predictions as action approximations.
-Diffusion models can also be used to extract the reward function Nuti et al.
-(2023) or value function Mazoure et al.
-(2023), and are prominent to estimate gradients with score-matching to solve offline optimization problems Suh et al.
-(2023).
-Enabling RL agents to generalize to multi-task and multi-agent scenarios remains a challenge due to their inherent complexity.
-Recent research has made progress in using diffusion models to improve the performance of policies in multi-task and multi-agent offline RL.
+(2023d) leverages consistency models as the policy class for fast training and sampling, while EDP Kang et al.
+(2023) achieves speed-up during training by using single-step model predictions as action approximations.
+Diffusion models are also used as the value function Mazoure et al.
+(2023) and representation model Venkatraman et al.
+(2023) to facilitate training of normal RL policies.
+Recent research has made progress in using diffusion models to improve the performance of policies in multitask and multi-agent offline RL.
 Multitask offline RL.
-Diffusion model is verified to have the potential to address the challenge of multi-task generalization in RL.
+Diffusion models are verified to have the potential to address the challenge of multi-task generalization in offline RL.
 He et al.
@@
 (2023a) leverages a hierarchical structure to achieve long-horizon multi-task control.
-Ni et al.
-(2023) and Liang et al.
-(2023) extend the idea of Diffuser Janner et al.
-(2022) into more specific sittings.
 MetaDiffuser Ni et al.
@@
 AdaptDiffuser Liang et al.
-(2023) combines bootstrapping and diffusion-based generative modeling together to enable the model to adapt to unseen tasks.
+(2023a) combines bootstrapping and diffusion-based generative modeling together to enable the model to adapt to unseen tasks.
 Multi-agent offline RL.
-Employing diffusion models to multi-agent RL helps model discrepant behaviors among agents and reduces approximation error.
+Using diffusion models in multi-agent RL helps model discrepant behaviors among agents and reduces approximation error.
 MADiff Zhu et al.
@@
 (2023c) incorporates the diffusion model into the policy classes to enhance learning and makes it possible to generalize to shifted environments well.
+5.2 Online RL
+Recently, there have been some works showing that diffusion models can also boost online RL training.
+Value estimations in online RL are noisier and change with the current policy, which poses additional challenges on training a multistep diffusion model.
+DIPO Yang et al.
+(2023a) proposes an action relabeling strategy to perform policy improvement at the data level, bypassing the potentially unstable value-guided training.
+Actions in the online rollout dataset are updated with gradient ascent, and the diffusion training objective is just supervised learning on the relabeled dataset.
+Chen et al.
+(2023d) conduct experiments to verify that consistency models with one-step sampling can naturally serve as online RL policies and achieve a balance between exploration and exploitation.
+Instead of using diffusion models as policies, Rigter et al.
+(2023) build a diffusion dynamic model to generate synthetic trajectories that are consistent with online RL policies.
+Applications of diffusion models in online RL are less explored compared to offline settings and merit further investigation.
 5.3 Imitation Learning
@@
 (2023).
-Diffusion models are already applied to goal-conditioned RL: Reuss et al.
+Diffusion models are already applied to goal-conditioned IL: Reuss et al.
 (2023) use a decoupled score-based diffusion model to learn an expressive goal-conditional policy.
@@
 (2023) build a unified diffusion policy to solve both goal-directed navigation and goal-agnostic exploration problems.
+Liang et al.
+(2023b) adopt a hierarchical structure where the high-level skills are determined by the current visual observation and language instructions.
+Therefore the low-level skill-conditioned planner can satisfy the user-specified multitask instructions.
 5.4 Trajectory Generation
-Trajectory generation aims to derive a dynamically feasible path that satisfies a set of constraints.
-In particular, we focus on generating human pose and robot interaction sequences, which are more related to the decision-making scenario.
+Trajectory generation aims to produce a dynamically feasible path that satisfies a set of constraints.
+We focus on using diffusion models to generate human pose and robot interaction sequences, which are more related to the decision-making scenario.
 Many works Zhang et al.
@@
 (2023a) harness diffusion models to synthesize a set of future frames depicting its planned actions in the future, after which control actions are extracted from the generated video.
-This approach makes it possible to train policies solely based on RGB videos and deploy learned policies to various robotic tasks Black et al.
+This approach makes it possible to train policies solely on RGB videos and deploy learned policies to various robotic tasks with varying dynamics Black et al.
 (2023); Gao et al.
@@
 UniSim Yang et al.
-(2023) uses diffusion models to build a universal simulator of real-world interaction by learning through combined diverse datasets.
-It can be utilized to train both high-level vision-language planners and low-level RL policies, demonstrating powerful emulation ability.
+(2023b) uses diffusion models to build a universal simulator of real-world interaction by learning through combined diverse datasets.
+It can be used to train both high-level vision-language planners and low-level RL policies, demonstrating powerful emulation ability.
 5.5 Data Augmentation
-Diffusion models have already been verified to be useful for data augmentation in the RL domain.
 Since diffusion models perform well in learning over multimodal or even noisy distributions, they can model original data distribution precisely.
@@
 (2023a) generate complete transitions of trained tasks via a diffusion model.
-It proves that such augmentation brings about significant policy improvement for both online and offline RL.
-6 Limitations and Remarks
-In this section, we list three limitations when applying diffusion models in RL, and include remarks on comparing diffusion-based generative modeling in RL to transformer-based autoregressive approaches.
-Application in online RL.
-Though diffusion models have contributed to offline RL from various perspectives, the dynamic and evolving nature of online RL introduces a significant challenge, as the data distribution sampled by the current policy can change over time.
-The primary concern is that effectively adapting diffusion models to changing data distributions requires a substantial amount of new data Lu et al.
-(2023b).
-These models are typically trained on fixed datasets, and incorporating enough data to ensure they can generalize and remain effective in dynamic real-world scenarios is a resource-intensive task.
-Balancing the need for adaptability with the requirement for extensive data is the primary consideration when applying diffusion models in online RL.
-It is promising to solve this dilemma with more lightweight diffusion models which can keep consistency as the data distribution changes during online interactions.
-Iterative sampling cost.
-The unique stepwise denoising mechanism of diffusion models makes the sampling procedure need to infer the trained model multiple times, significantly increasing the sampling cost.
-Although we can perform sampling acceleration techniques such as DDIM or DPM-Solver, the expensive inference cost limits the model to conduct high-frequency output in online interactions with virtual environments and real-world continuous control scenarios.
-The problem becomes more severe in those methods that generate a long trajectory but only take the first state or action for execution, such as Diffuser Janner et al.
-(2022) or Decision Diffuser Ajay et al.
-(2023).
-Chen et al.
-(2023d) incorporate the recently proposed Consistency Model Song et al.
-(2023) to enable sampling with one or two diffusion steps, and the performances are comparable to that of DDPM or DDIM with 50 steps.
-Variance in stochastic sampling.
-In traditional RL algorithms, a continuous control policy is represented by a state-conditioned Gaussian Haarnoja et al.
-(2018); Schulman et al.
-(2017).
-We can take its mean as action when deterministic execution is required.
-However, such deterministic policy is not possible when using diffusion models as the policy class.
-The randomness of diffusion sampling comes from both initial noise and per-step stochastic denoising.
-Although per-step stochasticity can be avoided if sampling with ODE-based methods like DDIM, the randomness in initial noisy samples remains inevitable.
-The high-variance policies can have a negative impact in environments with high accuracy or safety requirements.
-Existing works in RL rarely discuss this limitation, and sampling methods with reduced variances are expected.
-Comparison to transformer-based methods.
-The idea of using diffusion models commonly abstracts offline RL as a conditional generative modeling problem, differing from traditional RL approaches, which need value function approximations or policy gradient calculation.
-This supervised learning paradigm is similar to Decision Transformer (DT) Chen et al.
-(2021), a sequential modeling framework for RL tasks with transformer architecture.
-The essence of both approaches is to take advantage of the high expressiveness model, where diffusion models enjoy a strong distribution fitting ability to produce multi-modal, diverse, and accurate outputs, and transformers are adept in long-horizon sequence modeling and time correlation understanding.
-Qualitatively, this difference in expertise makes the diffusion-based approach more suitable for learning complex multi-modal tasks, and Transformer-based approaches Meng et al.
-(2021); Wen et al.
-(2022) are more preferred in correlated sequence modeling in time or agent (in multi-agent tasks) dimension.
+Lu et al.
+(2023b) directly train the diffusion model from the offline dataset or the online replay buffer and then generate samples for policy improvement.
+Analysis shows that both diversity and accuracy of data generated by diffusion models are higher than those generated by prior data augmentation methods.
+He et al.
+(2023a) deploy diffusion synthesizer on multi-task offline datasets and achieve better performance than that on single-task datasets.
+They claim that fitting on multiple tasks may enable implicit knowledge sharing across tasks, which also benefits from the multi-modal property of diffusion models.
+These works demonstrate that diffusion augmentation can bring significant improvement in policy learning for IL, online RL, and offline RL.
 6 Summary and Future Prospects
 This survey offers a comprehensive overview of contemporary research endeavors concerning the application of diffusion models in the realm of RL.
-According to the roles played by diffusion models, we categorize existing methodologies into using diffusion models as planners, policies, data synthesizers, and less popular roles such as value functions, latent representation models, etc.
+According to the roles played by diffusion models, we categorize existing methods into using diffusion models as planners, policies, data synthesizers, and less popular roles such as value functions, representation models, etc.
 By comparing each class of methods to traditional solutions, we can see how the diffusion model addresses some of the longstanding challenges in RL, i.e., restricted expressiveness, data scarcity, compounding error, and multitask generalization.
-Notwithstanding these merits, it is imperative to acknowledge the existence of non-negligible limitations in using diffusion models in RL due to some inherent properties in the training and sampling of diffusion models.
 It is worth emphasizing that the incorporation of diffusion models into RL remains an emerging field, and there are many research topics worth exploring.
-Here, we outline three prospective research directions, namely, retrieval-enhanced generation, integrating safety constraints, and composing different skills.
+Here, we outline four prospective research directions, namely, generative simulation, integrating safety constraints, retrieval-augmented generation, and composing different skills.
+Generative simulation.
+As shown in Fig.
+1, existing works use diffusion models to overcome certain limitations of previous solutions in both the agent and buffer parts.
+However, there has been a scarcity of research focused on using diffusion modeling to improve the environment.
+Gen2Sim Katara et al.
+(2023) uses text-to-image diffusion models to generate diverse objects in simulation environments, where RL policies are trained to learn robot manipulation skills.
+Besides generating objects in the scene, diffusion models have the potential for broader applications in generative simulation, such as the generation of various possible dynamics functions, reward functions, or opponents in multi-agent learning.
+Integrating safety constraints.
+Making decisions in real tasks often necessitates compliance with various safety constraints.
+Several safe RL methods transform a constrained RL problem to its unconstrained equivalent Achiam et al.
+(2017), which is then solved by generic RL algorithms.
+Policies acquired through these methods remain tailored to the specific constraint threshold specified during training.
+A recent research Liu et al.
+(2023) has extended the applicability of decision transformers to the context of safety-constrained settings, thereby enabling a single model to adapt to diverse thresholds by adjusting the input cost-to-go.
+Similarly, diffusion models have the potential to be deployed in safe RL by viewing safety constraints as sampling conditions.
+Ajay et al.
+(2023) demonstrate that a diffusion-based planner can combine different movement skills to produce new behaviors.
+Also, classifier-guided sampling can include new conditions by learning additional classifiers, while the parameters of the diffusion model remain unchanged Dhariwal and Nichol (2021).
+This makes the diffusion model promising for scenarios with new safety requirements after model training.
 Retrieval-enhanced generation.
@@
 Also, if the retrieval dataset is constantly updated, diffusion models have the potential to generate new behaviors without retraining.
-Integrating safety constraints.
-Utilizing RL models for practical applications often necessitates compliance with various safety constraints García and Fernández (2015); Gu et al.
-(2022).
-Several safe RL methods transform a constrained RL problem to its unconstrained equivalent Achiam et al.
-(2017); Tessler et al.
-(2018); Sootla et al.
-(2022), which is then solved by generic RL algorithms.
-Policies acquired through these methods remain tailored to the specific constraint threshold utilized during training.
-Recent research Liu et al.
-(2023); Zhang et al.
-(2023c) has extended the applicability of decision transformers to the context of safety-constrained settings, thereby enabling a single model to adapt to diverse thresholds by adjusting the input cost-to-go.
-Similarly, diffusion models exhibit substantial potential for deployment in the domain of safe RL.
-Ajay et al.
-(2023) demonstrate that a diffusion-based planner can combine different movement skills to produce new behaviors.
-In addition, classifier-guided sampling can add new conditions to generated samples simply by learning additional classifiers, while the parameters of the diffusion model itself remain unchanged Dhariwal and Nichol (2021).
-This feature makes the diffusion model promising to be effective in scenarios where new safety constraints can be included after model training.
 Composing different skills.
-Most current works deploy the generation ability of diffusion models on the raw state and action spaces.
-From the perspective of skill-based RL Shi et al.
-(2022); Nam et al.
-(2022), it is promising to break down complex tasks into smaller, more manageable sub-skills.
+From the perspective of skill-based RL, it is promising to break down complex tasks into smaller, more manageable sub-skills.
 Diffusion models excel in modeling multi-modal distributions, and since multiple sub-skills can be viewed as distinct modes within the distribution of possible behaviors, they offer a natural fit for this task.
@@
 Experiments in offline RL also suggest that diffusion models can share knowledge across skills and combine them up Ajay et al.
-(2023); He et al.
-(2023a), thus having the potential for zero-shot adaptation or continuous RL by composing different skills.
+(2023), thus having the potential for zero-shot adaptation or continuous RL by composing different skills.
 Acknowledgments